### v1

# Haptic bilateral teleoperation system for free-hand dental procedures

Abstract
Free-hand dental procedures are typically repetitive, time-consuming and require high precision and manual dexterity. Dental robots can play a key role in improving procedural accuracy and safety, enhancing patient comfort, and reducing operator workload. However, robotic solutions for free-hand procedures remain limited or completely lacking, and their acceptance is still low. To address this gap, we develop a haptic bilateral teleoperation system (HBTS) for free-hand dental procedures. The system includes a dedicated mechanical end-effector, compatible with standard clinical tools, and equipped with an endoscopic camera for improved visibility of the intervention site. By ensuring motion and force correspondence between the operator’s actions and the robot’s movements, monitored through visual feedback, we enhance the operator’s sensory awareness and motor accuracy. Furthermore, recognizing the need to ensure procedural safety, we limit interaction forces by scaling the motion references provided to the admittance controller based solely on measured contact forces. This ensures effective force limitation in all contact states without requiring prior knowledge of the environment. The proposed HBTS is validated in a dental scaling procedure using a dental phantom. The results show that the system improves the naturalness, safety, and accuracy of teleoperation, highlighting its potential to enhance free-hand dental procedures.
Index Terms:
Haptics, teleoperation, interaction control, free-hand dental procedures, eye-hand coordination
publicationid: pubid:
I Introduction
I-A Background
Robotics is rewriting the future of healthcare, emerging as a disruptive technology capable of optimizing and revolutionizing the way some medical procedures are conducted. Technological advancements enable robots to play an increasingly prominent role in the execution of complex medical procedures, where the required scale, accuracy, and precision are often beyond human capabilities. The successful application of such technology in the medical field is feeding a wave of innovation in dentistry, where robotics is changing the game [1].
Figure 1: Manual vs teleoperated dental procedures (pictorial views). In the upper right corner, is the console of The Senhance® Surgical System [2].
The World Health Organization estimates that oral diseases affect nearly 3.5 billion people worldwide. However, access to primary oral health services is limited due to the high costs of procedures, the unequal distribution of oral health professionals, and the lack of adequate healthcare facilities [3]. Furthermore, many procedures are conducted free-hand by uninformed operators, resulting in highly inefficient and disastrous outcomes.
Traditionally, routine, prosthetic, periodontal, and some surgical dental procedures are performed free-hand. Routine procedures include the treatment of dental caries and endodontic therapy, characterized by the mechanical excavation of the damaged or infected dental structure, typically utilizing burs attached to a rotary dental handpiece, followed by the replacement of the lost structure with a synthetic filling material [4]. Prosthetic procedures, on the other hand, focus on the replacement of missing or severely damaged teeth through the use of prostheses, such as inlays, crowns, or bridges. These procedures involve the preparation of the tooth, which entails the removal of hard tissue from the damaged or healthy tooth using rotary instruments mounted on dental handpieces [5]. Periodontal procedures include scaling and root planing (SRP), which involve the removal of plaque and calculus from the surfaces of the teeth and the smoothing of root surfaces, typically employing manual hand instruments and inserts mounted on sonic or ultrasonic handpieces [6]. Finally, although computer-assisted implant surgery (s-CAIS), dynamic computer-assisted implant surgery (d-CAIS), and robotic computer-assisted implant surgery (r-CAIS) have shown a significant increase in the accuracy of dental implant placement [7], many operators still prefer the free-hand method for dental implantology [8].
Each of these procedures, even the simplest one, requires accuracy, manual skills and dexterity on the part of the operator. Indeed, the latter has to cope with a highly unstructured environment, which is full of natural obstacles that occlude the view of the surgical site and reduce the space for maneuvering, requiring constant changes in the orientation of the surgical tool and the operator’s posture. The successful outcome of a procedure, as well as the comfort for the patient, therefore, depend on both human factors, i.e., the experience and skills of the human operator [9], the adequacy of the instrumentation used and the duration of the treatments [10], and environmental factors, i.e., difficult anatomical conditions and less accessibility to the surgical site [11]. In particular, the latter factors can significantly affect the operator’s comfort. In [12] a high prevalence of musculoskeletal disorders among dentists is reported. The main causes of these disorders can be traced back to an incorrect application of ergonomics, such as awkward postures, repetitive and prolonged movements, use of small or narrow tool handles, exertion of forces on hard or sharp objects, and exposure to vibrations [13]. Therefore, in the process of training new operators, much emphasis is placed on learning skills related to space utilization and positioning of surgical and non-surgical instruments, e.g., mirrors, to overcome occlusions, while maintaining healthy and comfortable posture [14]. In addition to this, to ensure the safety and effectiveness of the procedure, new operators must acquire skills related to the correct amount of pressure to apply to the site of interest, depending on the type of tool used. Consequently, the sense of touch plays a key role in achieving successful procedures. This is made evident by emerging trends in teaching, which make use of virtual environments and haptic devices (HD)s for training new operators [15, 16].
In contrast to humans, robots are capable of achieving accuracy, precision, efficiency and scales beyond human capabilities. For this reason, the role of dental robots is to improve dental procedures, both in terms of patient satisfaction, i.e., broader access to care, reduced treatment time, successful outcome of the procedure, fast recovery, and in terms of operator comfort, i.e., reduced workload intensity, increased visibility of the surgical field, improved ergonomics [17], as shown in Figure 1. Haptic teleoperations are a technology that combines the robot’s skills with the flexibility, dexterity, and cognitive capabilities of humans. This technology is applied in multiple domains, such as telesurgery and space exploration, as it overcomes barriers in terms of distance and hazardous environments, and scales beyond human reach. By providing multiple types of sensory information, i.e. visual and haptic feedback, these systems enable the human operator to experience telepresence, defined as the perception of interacting with the remote environment as if it were directly encountered [18]. Furthermore, by integrating different haptic feedback and guidance modalities [19, 20], and shared control approaches [21], they contribute to reducing errors and enhance the efficiency, accuracy and safety of procedures, thus extending human capabilities, while also improving operator comfort and confidence.
I-B Related works
The majority of research in robotics for dentistry is focused on oral implantology [22]. Robotic oral implant surgery offers significant benefits in terms of implant accuracy in partially [23] or totally endetulous patients, because of the control of angular deviation and axial errors, also compared to a s-CAIS system [24]. Better placement accuracy, however, can be achieved at the expense of surgical efficiency, as shown in [25], which compares surgical efficiency and implant positioning accuracy among three types of robotic systems: passive, semi-active, and active robots. Such technologies also enable flapless dental implant placement, speeding up recovery time, reducing postoperative pain and swelling, and enhancing patient comfort. [26] shows how a telerobotic-assisted drilling system can enhance the accuracy of the human operator while preserving awareness of the procedure through an effective transmission of cutting forces and vibrations.
On the other hand, the state-of-the-art shows considerably less research interest in other free-hand procedures [1]. In the context of prosthetic procedures, robotics can enhance the precision of manual grinding using a turbine drill during tooth preparation [27] and can also compensate for hand tremors and the lack of tactile feedback in laser ablation [28, 29], improving accuracy [30] and efficiency [31]. In endodontics, robotics can automate and enhance precision and efficiency in root canal treatments [32, 33], outperforming current approaches also in effectively addressing biological and mechanical traits associated with biofilm recalcitrance and reinfection [34]. In periodontal procedures, robotic systems have the potential to increase accuracy, reducing the risk of errors and trauma to tissues by optimizing applied force, as with other free-hand procedures. They can also improve the ergonomics for the operator, and comfort for the patient by speeding up recovery time. However, for such procedures robotic technologies are still lacking [35], and there is no evidence of such advantages.
Thus, although robotics emerges as a transformative technology in the context of dental practice, there are still many limitations to consider. First of all, the quality of research is still insufficient in terms of clinical validation and technology readiness level (TRL) [36]. This is heavily influenced by the high costs of these robotic systems and the need to ensure their compatibility with existing clinical tools [37]. Second, the most widely used category of robots is fully autonomous ones. However, the use of this type of robots requires the development of state-of-the-art algorithms and advanced capabilities in perception, recognition, and exception handling [37]. In this context, the presence of a human, who must possess digital and mechanical skills in addition to clinical expertise, is still necessary not only as a supervisor but also as an operator. Considering the absence of haptic feedback in most robotic systems, which would improve the efficiency and effectiveness of procedures and increase operator acceptance [38], as well as the need to provide access to care in areas with a shortage of professionals, haptic teleoperations represent a viable solution. These systems would enable the remote execution of procedures with the accuracy and efficiency of robots while preserving human cognitive abilities. Moreover, they would allow the acquisition of human control strategies directly exercised in the robot state space, simplifying the process of transferring the knowledge from the human to the robot [39] and enabling the effective and objective evaluation of operator skills during training [40]. Finally, research tends to prioritize oral implantology over other dental procedures, despite the high number of free-hand routine procedures performed daily and patient preferences, as they tend to accept these technologies more readily when the degree of invasiveness is lower [41]. Considering these limitations, [1] recommends to shift the focus from oral implantology procedures towards less invasive ones, which are better suited as forerunners in the field, and also to invest in educational robotics research in academia as a valuable tool to introduce dental robotics and overcome the barrier of acceptance of robotic systems among future dentists.
I-C Contributions
In our previous work, we proposed an HBTS to facilitate the execution of precision tasks and increase the safety of interaction using haptic teleoperations [42]. Motivated by the limitations or complete absence of robotic solutions for the most repetitive free-hand dental procedures, and recognizing the importance of haptic teleoperations in enhancing human capabilities while preserving dexterity, cognitive skills, and procedural awareness, in this paper we present an extended version of the system. Our contribution is an HBTS for free-hand dental procedures, including:
  • a dedicated mechanical end-effector, compatible with existing clinical tools and featuring an endoscopic camera, to improve visibility at the intervention site;
  • a virtualized force feedback, consistent with the interaction observed through the visual feedback of the eye-in-hand camera, which enhances the haptic experience for the operator;
  • a control system that ensures procedural safety, improves accuracy, and preserves manual dexterity, while providing motion correspondence between the operator’s movements and those of the end-effector, as observed in the camera’s view.
We validate the proposed HBTS on a free-hand procedure, consisting of dental scaling.
II Preliminaries
II-A Notation
We denote by ${\bf T}_{b}^{a}\in SE(3)$ the homogeneous transformation matrix from frame $a$ to frame $b$, in the form:
${\bf T}_{b}^{a}=\begin{pmatrix}{\bf R}_{b}^{a}&{\bf p}_{b}^{a}\\ {\bf 0}^{\top}&1\end{pmatrix},$ (1)
where ${\bf R}_{b}^{a}\in SO(3)$ is the rotation matrix from frame $a$ to frame $b$, and ${\bf p}_{b}^{a}\in\mathbb{R}^{3\times 1}$ represents $b$’s origin in frame $a$.
We let ${\bf x}_{b}^{a}$ be the pose of the frame $b$ with respect to the frame $a$, corresponding to ${\bf T}_{b}^{a}$, in the form:
${\bf x}_{b}^{a}=\begin{bmatrix}{\bf p}_{b}^{a}\\ {\bf q}_{b}^{a}\end{bmatrix},$ (2)
where ${\bf q}_{b}^{a}\in\mathbb{H}_{1}$ is the quaternion associated with ${\bf R}_{b}^{a}$.
We denote by ${\bf R}_{\bf a}^{\mathcal{F}}(\alpha)\in SO(3)$ the elementary rotation of an angle $\alpha$ about a generic axis $\bf a$ of a frame $\mathcal{F}$.
We denote by ${\bf v}^{a}_{b}\in\mathbb{R}^{6\times 1}$ the twist of the frame $b$ with respect to the frame $a$, in the form:
${\bf v}^{a}_{b}=\begin{bmatrix}\dot{\bf p}^{a}_{b}\\ {\boldsymbol{\omega}}^{a}_{b}\end{bmatrix},$ (3)
where $\dot{\bf p}^{a}_{b},{\boldsymbol{\omega}}^{a}_{b}$ are the linear and angular velocities, respectively.
II-B Haptic device and robot reference frames
The frames used by the HD and the robot are reported in Figure 2. With regard to the HD, two frames are introduced:
  • $\mathcal{H}_{b}$ is the base reference frame, in which the HD’s stylus poses and the feedback forces are expressed;
  • $\mathcal{H}_{e}$ is the HD’s stylus frame.
Regarding the robot, four frames are defined:
  • $\mathcal{R}_{b}$ is the base reference frame, in which motion references are expressed;
  • $\mathcal{R}_{e}$ is the robot’s end-effector frame (computed from forward kinematics of joint measurements), corresponding to the tool center point (TCP);
  • $\mathcal{R}_{ed}$ is the desired robot’s end-effector frame, input to admittance control;
  • $\mathcal{R}_{c}$ is the compliant frame, output by the admittance control law.
We note that, after admittance control, motion control guarantees that $\mathcal{R}_{e}$ tracks $\mathcal{R}_{c}$, as discussed next.
Figure 2: Haptic device and robot reference frames.
III Methodology
III-A System overview
The HBTS we propose, shown in Figure 3, is based on a position-force architecture [43], coupling the leader (LR) and follower (FR) robots. The LR is an impedance HD, sensing the pose of the human operator and rendering haptic feedback according to the remote interaction between the FR and the environment. The HD is controlled by a leader controller, featuring an eye-hand coordination controller (EHCC) and a haptic feedback controller (HFC). The former is responsible for capturing the HD’s stylus poses, filtering the human tremors, and generating the corresponding robot end-effector poses, so that the robot follows the stylus motion in coordination with the visual feedback. The latter is responsible for rendering the force feedback on the HD to let the human operator perceive the remote environment. Both controllers record data throughout the entire procedure. A detailed description of the objectives and operation of the two controllers is provided in Section III-B and Section III-D, respectively. The FR is a position-controlled robot, instructed to track the HD through an admittance controller, which provides the latter with a configurable degree of compliance on its target pose with respect to the environment. Such controller has three key features. First, it ensures the stability of the FR, even in case of loss of communication, since, in case of a lack of reference from the EHCC, the controller keeps following the last one received. Second, it avoids damage to the arm structure and manipulated environment, since it adapts the received references to ensure compliant interaction with the surroundings. Finally, it limits the interaction forces by scaling the received motion references through a free-hand force limitation strategy embedded in the controller itself. The force limitation strategy is explained in Section III-C2.
Figure 3: Haptic bilateral teleoperation system scheme.
III-B Eye-hand coordination controller
In an HBTS, visual feedback plays a crucial role in overcoming the limitations caused by distance. In teleoperated free-hand dental procedures, using an eye-in-hand camera is the most effective solution [44]. This approach enables the operator to maintain a consistently centered view of the operative field, offering a significant advantage over eye-to-hand camera configurations, where the view may be off-center with respect to the area of interest or obstructed by physical obstacles. In this context, the human operator is naturally guided by the video stream transmitted by the camera to provide commands to the FR via the HD. However, in the eye-in-hand configuration, a rotation of the tool (unavoidable, to preserve dexterity) causes a rotation of the viewing angle. This creates a mismatch between the movements expected by the operator and those actually executed by the robot in the camera view, compromising the accuracy of the task.
The da Vinci Surgical SystemTM, which requires the insertion of multiple tools to implement an eye-to-hand strategy, provides the motion correspondence between the operator’s movements and those of the end-effector, observed in the camera view, by projecting the endoscope image above the surgeon’s hands by means of mirrored overlay optics [45]. In our eye-hand coordination controller, we provide the motion correspondence at the command level, by creating a coupling between the HD’s stylus and the robot’s TCP, and appropriately converting the motion of the former into the motion of the FR. To enhance the naturalness of the teleoperation, a rotational coupling is also established between the HD’s stylus and the FR’s TCP. This includes defining the desired relative rotation ${\bf R}_{\mathcal{H}_{e}}^{\mathcal{R}_{e}}$, and performing manually the alignment. This is done by the human operator during what we indicate as engagement phase.
III-B1 Engagement phase
During the engagement phase, the EHCC retrieves the homogeneous transformation matrix of the HD’s stylus ${\bf T}_{\mathcal{H}_{e}}^{\mathcal{H}_{b}}(k)$, and converts it into a robot’s end-effector homogeneous transformation matrix with the following kinematic equations:
${\bf R}_{\mathcal{R}_{ed}}^{\mathcal{R}_{b}}(k)$ $={\bf R}_{\mathcal{H}_{b}}^{\mathcal{R}_{b}}{\bf R}_{\mathcal{H}_{e}}^{\mathcal{H}_{b}}(k)\left({\bf R}_{\mathcal{H}_{e}}^{\mathcal{R}_{e}}\right)^{\top},$ (4)
${\bf p}_{\mathcal{R}_{ed}}^{\mathcal{R}_{b}}(k)$ $={\bf p}_{\mathcal{R}_{e}}^{\mathcal{R}_{b}}(k),$ (5)
where ${\bf R}_{\mathcal{H}_{b}}^{\mathcal{R}_{b}}$ is the fixed rotation between $\mathcal{H}_{b}$ and $\mathcal{R}_{b}$. We note that the position of the HD’s stylus is not relevant for coupling, so it is converted to the current position of the robot’s TCP. The obtained frame $\mathcal{R}_{ed}$ is real-time displayed in a visualization tool, together with the actual robot end-effector one $\mathcal{R}_{e}$, as shown in Figure 4, so that the human operator, by acting on $\mathcal{H}_{e}$, can align the frame $\mathcal{R}_{ed}$ with the frame $\mathcal{R}_{e}$. Once the alignment is accomplished, i.e.  $\left[{\bf R}_{\mathcal{R}_{e}}^{\mathcal{R}_{b}}(k)\right]^{\top}{\bf R}_{\mathcal{R}_{ed}}^{\mathcal{R}_{b}}(k)={\bf I}_{3}$, where ${\bf I}_{3}\in SO(3)$ is the identity matrix, the engagement phase ends, i.e. HD and FR are aligned, and the human operator can begin teleoperation.
Figure 4: Engagement phase supported by the visualization tool.
III-B2 Human tremors filtering
At each control cycle, the EHCC retrieves the actual pose ${\bf x}_{\mathcal{H}_{e}}^{\mathcal{H}_{b}}(k)$ of the HD’s stylus, corresponding to ${\bf T}_{\mathcal{H}_{e}}^{\mathcal{H}_{b}}(k)$. Exploiting such a pose to generate a command for the robot would result in transferring the human hand motion tremors, further accentuated by the control frequency. Therefore, ${\bf x}_{\mathcal{H}_{e}}^{\mathcal{H}_{b}}(k)$ is filtered online by a moving average filter for poses. At the $k$-th control cycle, the filtered position $\overline{{\bf p}}_{\mathcal{H}_{e}}^{\mathcal{H}_{b}}(k)$ is obtained from the cumulative sum ${\bf s}(k)$ of samples within a window of fixed size $n$:
$\overline{{\bf p}}_{\mathcal{H}_{e}}^{\mathcal{H}_{b}}(k)=\frac{{\bf s}(k)}{m}$ (6)
where $m\leq n$ is the number of samples in the moving average window at the $k$-th control cycle, and the cumulative sum ${\bf s}(k)$ is:
${\bf s}(k)=\begin{cases}{\bf s}(k-1)+{\bf p}_{\mathcal{H}_{e}}^{\mathcal{H}_{b}}(k),&\text{if}~{}k<n;\\ {\bf s}(k-1)+{\bf p}_{\mathcal{H}_{e}}^{\mathcal{H}_{b}}(k)-{\bf p}_{\mathcal{H}_{e}}^{\mathcal{H}_{b}}(k-n),&\text{otherwise.}\end{cases}$ (7)
The filtered quaternion $\overline{\bf q}_{\mathcal{H}_{e}}^{\mathcal{H}_{b}}(k)$ is the eigenvector associated with the maximum eigenvalue of the ${\bf M}(k)$ matrix, of proper size [46], defined as:
${\bf M}(k)=\frac{{\bf c}(k)}{m},$ (8)
where ${\bf c}(k)$ is the cumulative sum of samples within a window of fixed size $n$, updated at each control cycle:
${\bf c}(k)=\begin{cases}{\bf c}(k-1)+{\bf q}_{\mathcal{H}_{e}}^{\mathcal{H}_{b}}(k)\left[{\bf q}_{\mathcal{H}_{e}}^{\mathcal{H}_{b}}(k)\right]^{\top},&\text{if}~{}k<n;\\ {\bf c}(k-1)+{\bf q}_{\mathcal{H}_{e}}^{\mathcal{H}_{b}}(k)\left[{\bf q}_{\mathcal{H}_{e}}^{\mathcal{H}_{b}}(k)\right]^{\top}+&\\ \quad-{\bf q}_{\mathcal{H}_{e}}^{\mathcal{H}_{b}}(k-n)\left[{\bf q}_{\mathcal{H}_{e}}^{\mathcal{H}_{b}}(k-n)\right]^{\top},&\text{otherwise.}\end{cases}$ (9)
Keeping the cumulative sums ${\bf s}(k)$ and ${\bf c}(k)$ enables achieving $\mathcal{O}(1)$ computational complexity.
III-B3 Eye-hand coordination
In an HBTS, coupling two robots with different kinematics requires the use of mapping and scaling techniques. The literature shows that this coupling can be achieved by converting the motion of the HD into that of the FR through simple kinematic transformations, and appropriately scaling it to either expand the workspace size of the HD or enhance control accuracy [47]. In the eye-in-hand configuration, converting quantities between the two spaces without considering the viewing angle would lead to a loss of motion correspondence between the movements expected by the operator and those actually executed by the robot in the camera view. Therefore, in our EHCC, we transfer the HD’s twist to the FR through an eye-hand coordination mapping and introduce dynamic scaling factors to improve control accuracy based on the task requirements.
Let $\Delta\overline{{\bf p}}_{\mathcal{H}_{e}}^{\mathcal{H}_{b}}(k)$ be the HD’s displacement:
$\Delta\overline{{\bf p}}_{\mathcal{H}_{e}}^{\mathcal{H}_{b}}(k)=\overline{{\bf p}}_{\mathcal{H}_{e}}^{\mathcal{H}_{b}}(k)-\overline{{\bf p}}_{\mathcal{H}_{e}}^{\mathcal{H}_{b}}(k-1),$ (10)
and $\overline{{\bf R}}^{\mathcal{H}_{e}(k-1)}_{\mathcal{H}_{e}(k)}$ the relative rotation:
$\overline{{\bf R}}^{\mathcal{H}_{e}(k-1)}_{\mathcal{H}_{e}(k)}=\begin{bmatrix}r_{11}&r_{12}&r_{13}\\ r_{21}&r_{22}&r_{23}\\ r_{31}&r_{32}&r_{33}\end{bmatrix},$ (11)
obtained from the quaternion $\overline{\bf q}_{\mathcal{H}_{e}(k)}^{\mathcal{H}_{e}(k-1)}$:
$\overline{\bf q}_{\mathcal{H}_{e}(k)}^{\mathcal{H}_{e}(k-1)}=\left[\overline{\bf q}_{\mathcal{H}_{e}}^{\mathcal{H}_{b}}(k-1)\right]^{\top}\overline{\bf q}_{\mathcal{H}_{e}}^{\mathcal{H}_{b}}(k).$ (12)
At each control cycle, the EHCC computes the HD’s stylus twist:
${\bf v}_{\mathcal{H}_{e}}^{\mathcal{H}_{b}}(k)=\begin{bmatrix}\dfrac{\Delta\overline{{\bf p}}_{\mathcal{H}_{e}}^{\mathcal{H}_{b}}(k)}{T_{s}}\\[8.5359pt] \dfrac{\vartheta{\bf r}}{T_{s}}\end{bmatrix},$ (13)
where $T_{s}$ is the control period and ${\bf r}$ and $\vartheta$ are the axis and the angle corresponding to $\overline{{\bf R}}^{\mathcal{H}_{e}(k-1)}_{\mathcal{H}_{e}(k)}$, respectively:
$\vartheta$ $=\cos^{-1}{\left(\dfrac{r_{11}+r_{22}+r_{33}-1}{2}\right)},$ (14)
${\bf r}$ $=\frac{1}{2\sin{\vartheta}}\begin{bmatrix}r_{32}-r_{23}\\ r_{13}-r_{31}\\ r_{21}-r_{12}\end{bmatrix}.$ (15)
From (13), the corresponding desired FR’s end-effector twist is obtained as:
$\begin{split}{\bf v}_{\mathcal{R}_{ed}}^{\mathcal{R}_{b}}(k)=&\left[{\bf I}_{2}\otimes{\bf R}_{\mathcal{R}_{e}}^{\mathcal{R}_{b}}(k)\right]{\boldsymbol{\Lambda}}\left[{\bf I}_{2}\otimes{\bf R}_{\mathcal{H}_{e}}^{\mathcal{R}_{e}}(k)\right]\\ &\cdot\left[{\bf I}_{2}\otimes{\bf R}_{\mathcal{H}_{e}}^{\mathcal{H}_{b}}(k)\right]^{\top}{\bf v}_{\mathcal{H}_{e}}^{\mathcal{H}_{b}}(k).\end{split}$ (16)
where $\otimes$ denotes the Kronecker product, ${\boldsymbol{\Lambda}}$ is the scaling diagonal matrix, of proper size, and ${\bf R}_{\mathcal{H}_{e}}^{\mathcal{R}_{e}}(k)$ is:
${\bf R}_{\mathcal{H}_{e}}^{\mathcal{R}_{e}}(k)={\bf R}_{\mathcal{H}_{e}}^{\mathcal{R}_{e}}{\bf R}_{\bf z}^{\mathcal{H}_{e}}(\varphi(k)).$ (17)
We note that ${\bf R}_{\mathcal{H}_{e}}^{\mathcal{R}_{e}}$ is the fixed rotation defining the mapping between quantities in $\mathcal{H}_{e}$ and $\mathcal{R}_{e}$, and $\bf z$ is the axis of $\mathcal{H}_{e}$ mapped to the ${\bf o}$ axis of $\mathcal{R}_{e}$, according to ${\bf R}_{\mathcal{H}_{e}}^{\mathcal{R}_{e}}$. The ${\bf o}$ axis represents the optical axis of the camera in $\mathcal{R}_{e}$, which we assume to coincide with the ${\bf z}$ axis of $\mathcal{R}_{e}$ for negligible values of the camera tilt angle. Therefore, a rotation of $\varphi(k)$ about the ${\bf z}$ axis of $\mathcal{H}_{e}$ results in a viewing angle of $\varphi(k)$, as shown in Figure 5.
[…]
Figure 5: Visualization of the rotational coupling between $\mathcal{H}_{e}$ and $\mathcal{R}_{e}$.
The rotation matrix ${\bf R}_{\bf z}^{\mathcal{H}_{e}}(\varphi(k))$ ensures the motion correspondence between the operator’s movements and those of the FR’s TCP in the camera view, by appropriately rotating the desired velocities according to the viewing angle. By setting ${\bf R}_{\bf z}^{\mathcal{H}_{e}}(\varphi(k))={\bf I}_{3}$ the mapping between twists becomes viewing angle independent:
${\bf R}_{\mathcal{H}_{e}}^{\mathcal{R}_{e}}(k)={\bf R}_{\mathcal{H}_{e}}^{\mathcal{R}_{e}},\forall k.$ (18)
The desired FR’s TCP homogeneous transformation ${\bf T}_{\mathcal{R}_{ed}}^{\mathcal{R}_{b}}(k)$ is obtained by numerically integrating the desired twist in (16). As for the desired rotation ${\bf R}_{\mathcal{R}_{ed}}^{\mathcal{R}_{b}}(k)$, it is obtained as:
${\bf R}_{\mathcal{R}_{ed}}^{\mathcal{R}_{b}}(k)={\bf R}_{\mathcal{R}_{e}}^{\mathcal{R}_{b}}(k-1){\bf R}_{\mathcal{R}_{ed}(k)}^{\mathcal{R}_{e}(k-1)},$ (19)
where ${\bf R}_{\mathcal{R}_{ed}(k)}^{\mathcal{R}_{e}(k-1)}$ is the rotation matrix corresponding to the axis $\frac{{\bf r}}{\left\lVert{\bf r}\right\rVert}$ and angle $\vartheta$:
${\bf r}$ $=\left({\bf R}_{\mathcal{R}_{e}}^{\mathcal{R}_{b}}(k-1)\right)^{\top}{\boldsymbol{\omega}}_{\mathcal{R}_{ed}}^{\mathcal{R}_{b}}(k)T_{s},$ (20)
$\vartheta$ $=\left\lVert{\bf r}\right\rVert.$ (21)
As for the desired position ${\bf p}_{\mathcal{R}_{ed}}^{\mathcal{R}_{b}}(k)$, it is obtained as:
${\bf p}_{\mathcal{R}_{ed}}^{\mathcal{R}_{b}}(k)={\bf p}_{\mathcal{R}_{ed}}^{\mathcal{R}_{b}}(k-1)+\dot{\bf p}_{\mathcal{R}_{ed}}^{\mathcal{R}_{b}}(k)T_{s}.$ (22)
From ${\bf T}_{\mathcal{R}_{ed}}^{\mathcal{R}_{b}}(k)$ the desired FR’s TCP pose is obtained as:
${\bf x}_{\mathcal{R}_{ed}}^{\mathcal{R}_{b}}(k)=\begin{bmatrix}{\bf p}_{\mathcal{R}_{ed}}^{\mathcal{R}_{b}}(k)^{\top}&{\bf q}_{\mathcal{R}_{ed}}^{\mathcal{R}_{b}}(k)^{\top}\end{bmatrix}^{\top}.$ (23)
The eye-hand coordination mapping and scaling can be turned into an algorithmic procedure with steps provided in Algorithm 1.
Algorithm 1 Eye-hand coordination mapping and scaling
1:The HD’s actual $\overline{\bf x}_{\mathcal{H}_{e}}^{\mathcal{H}_{b}}(k)$ and previous $\overline{\bf x}_{\mathcal{H}_{e}}^{\mathcal{H}_{b}}(k-1)$ filtered poses, and the previous desired pose ${\bf x}_{\mathcal{R}_{ed}}^{\mathcal{R}_{b}}(k-1)$
2:The desired FR’s TCP pose ${\bf x}_{\mathcal{R}_{ed}}^{\mathcal{R}_{b}}(k)$
3:Compute ${\bf v}_{\mathcal{H}_{e}}^{\mathcal{H}_{b}}(k)$ using (13)
4:Compute ${\bf v}_{\mathcal{R}_{ed}}^{\mathcal{R}_{b}}(k)$ using (16):
  1. 1.
    Convert ${\bf v}_{\mathcal{H}_{e}}^{\mathcal{H}_{b}}(k)$ into ${\bf v}_{\mathcal{R}_{ed}}^{\mathcal{R}_{e}}(k)$ and rotate it according to the viewing angle using (17):
    $\begin{split}{\bf v}_{\mathcal{R}_{ed}}^{\mathcal{R}_{e}}(k):=&\left[{\bf I}_{2}\otimes{\bf R}_{\mathcal{H}_{e}}^{\mathcal{R}_{e}}(k)\right]\left[{\bf I}_{2}\otimes{\bf R}_{\mathcal{H}_{e}}^{\mathcal{H}_{b}}(k)\right]^{\top}{\bf v}_{\mathcal{H}_{e}}^{\mathcal{H}_{b}}(k)\end{split}$
  2. 2.
    Scale ${\bf v}_{\mathcal{R}_{ed}}^{\mathcal{R}_{e}}(k)$ using $\boldsymbol{\Lambda}$ and convert it into ${\bf v}_{\mathcal{R}_{ed}}^{\mathcal{R}_{b}}(k)$:
    ${\bf v}_{\mathcal{R}_{ed}}^{\mathcal{R}_{b}}(k)=\left[{\bf I}_{2}\otimes{\bf R}_{\mathcal{R}_{e}}^{\mathcal{R}_{b}}(k)\right]{\boldsymbol{\Lambda}}{\bf v}_{\mathcal{R}_{ed}}^{\mathcal{R}_{e}}(k)$
5:Compute ${\bf T}_{\mathcal{R}_{ed}}^{\mathcal{R}_{b}}(k)$ from ${\bf v}_{\mathcal{R}_{ed}}^{\mathcal{R}_{b}}(k)$ using (19) and (22)
6:Extract ${\bf x}_{\mathcal{R}_{ed}}^{\mathcal{R}_{b}}(k)$ from ${\bf T}_{\mathcal{R}_{ed}}^{\mathcal{R}_{b}}(k)$
III-C Robot interaction control
III-C1 Admittance model
Let ${\bf x}_{\mathcal{R}_{ed}}^{\mathcal{R}_{b}}(k)=\begin{bmatrix}{\bf p}_{\mathcal{R}_{ed}}^{\mathcal{R}_{b}}(k)^{\top}&{\bf q}_{\mathcal{R}_{ed}}^{\mathcal{R}_{b}}(k)^{\top}\end{bmatrix}^{\top}$ be the desired pose and $\mathcal{R}_{c}$ the compliant frame specified by the quantities ${\bf p}_{\mathcal{R}_{c}}^{\mathcal{R}_{e}}(k)$, $\dot{\bf p}_{\mathcal{R}_{c}}^{\mathcal{R}_{e}}(k)$, $\ddot{\bf p}_{\mathcal{R}_{c}}^{\mathcal{R}_{e}}(k)$ and ${\bf q}_{\mathcal{R}_{c}}^{\mathcal{R}_{e}}(k)$. We do not consider compliance along the rotational axes, therefore, we assume ${\bf q}_{\mathcal{R}_{c}}^{\mathcal{R}_{e}}(k)={\bf q}_{\mathcal{R}_{ed}}^{\mathcal{R}_{e}}(k)$. When the FR interacts with the environment, $\mathcal{R}_{c}$ can be completely defined from ${\bf p}_{\mathcal{R}_{ed}}^{\mathcal{R}_{e}}(k)$, $\dot{\bf p}_{\mathcal{R}_{ed}}^{\mathcal{R}_{e}}(k)$ and $\ddot{\bf p}_{\mathcal{R}_{ed}}^{\mathcal{R}_{e}}(k)$, and the contact forces ${\bf f}^{\mathcal{R}_{e}}(k)\in\mathbb{R}^{3\times 1}$, by integrating the mechanical impedance equation:
${\bf M}_{d}\ddot{\tilde{{\bf p}}}^{\mathcal{R}_{e}}(k)+{\bf K}_{D}\dot{\tilde{{\bf p}}}^{\mathcal{R}_{e}}(k)+{\bf K}_{P}\tilde{{\bf p}}^{\mathcal{R}_{e}}(k)=-\tilde{\bf f}^{\mathcal{R}_{e}}(k),$ (24)
where ${\bf M}_{d},{\bf K}_{D},{\bf K}_{P}$ are the mass, damping, and stiffness diagonal matrices, respectively, of proper size, which are used to impose specific second-order mass-spring-damper dynamics; $\tilde{\bf p}^{\mathcal{R}_{e}}(k)={\bf p}^{\mathcal{R}_{e}}_{\mathcal{R}_{ed}}(k)-{\bf p}^{\mathcal{R}_{e}}_{\mathcal{R}_{c}}(k)$ is the position error and $\dot{\tilde{\mathbf{p}}}^{\mathcal{R}_{e}}(k)$ and $\ddot{\tilde{\mathbf{p}}}^{\mathcal{R}_{e}}(k)$ its time derivatives; $\tilde{{\bf f}}^{\mathcal{R}_{e}}(k)={\bf f}_{d}^{\mathcal{R}_{e}}(k)-{\bf f}_{e}^{\mathcal{R}_{e}}(k)$ is the force tracking error, defined as the difference between the desired force ${\bf f}_{d}^{\mathcal{R}_{e}}(k)$ and the exerted force ${\bf f}_{e}^{\mathcal{R}_{e}}(k)$, which is equal and opposite to the measured force. We take ${\bf f}_{d}^{\mathcal{R}_{e}}(k)={\bf 0},\forall k\in\mathbb{N}_{0}$ to make the robot behave compliantly with the environment and $\dot{\bf p}^{\mathcal{R}_{e}}_{\mathcal{R}_{ed}}(k)=\ddot{\bf p}^{\mathcal{R}_{e}}_{\mathcal{R}_{ed}}(k)={\bf 0},\forall k\in\mathbb{N}_{0}$ to reach a complete stop of the FR’s end-effector upon reaching the target position, effectively preventing any collisions caused by further movement. From (24), the compliant position reference to be commanded to the robot is obtained as:
${\bf p}_{\mathcal{R}_{c}}^{\mathcal{R}_{b}}(k)={\bf p}_{\mathcal{R}_{e}}^{\mathcal{R}_{b}}(k)+{\bf R}_{\mathcal{R}_{e}}^{\mathcal{R}_{b}}(k){\bf p}_{\mathcal{R}_{c}}^{\mathcal{R}_{e}}(k).$ (25)
III-C2 Free-hand force limitation strategy
Assuming that the environment can be approximated by a linear spring model, the exerted force at the steady state [48] is:
${\bf f}_{ss}^{\mathcal{R}_{e}}={\bf K}_{eq}\left({\bf p}_{\mathcal{E}_{r}}^{\mathcal{R}_{e}}-{\bf p}_{\mathcal{R}_{ed}}^{\mathcal{R}_{e}}\right),$ (26)
where ${\bf K}_{eq}=({\bf K}_{P}{\bf K}_{e})({\bf K}_{P}+{\bf K}_{e})^{-1}$ is the equivalent stiffness of the admittance stiffness ${\bf K}_{P}$ and environmental stiffness ${\bf K}_{e}$ [49], and ${\bf p}_{\mathcal{E}_{r}}^{\mathcal{R}_{e}}$ is the environment rest position. From (26), it follows that interaction forces can be limited through a suitable reference ${\bf p}_{\mathcal{R}_{ed}}^{\mathcal{R}_{e}}$ provided to the controller. Accurate limitation requires perfect knowledge of the environment that is particularly challenging to achieve, especially in a medical environment, in which the FR interacts with objects of varying stiffness and non-uniform geometry. Therefore, in the proposed strategy, the position reference:
${\bf p}_{\mathcal{R}_{ed}}^{\mathcal{R}_{e}}(k)=\left[{\bf R}_{\mathcal{R}_{e}}^{\mathcal{R}_{b}}(k)\right]^{\top}{\bf p}_{\mathcal{R}_{ed}}^{\mathcal{R}_{b}}(k)-\left[{\bf R}_{\mathcal{R}_{e}}^{\mathcal{R}_{b}}(k)\right]^{\top}{\bf p}_{\mathcal{R}_{e}}^{\mathcal{R}_{b}}(k),$ (27)
is dynamically scaled based solely on the contact forces:
$\underline{{\bf p}}_{\mathcal{R}_{ed}}^{\mathcal{R}_{e}}(k)=\frac{1}{1+K_{s}\left\lVert{\bf f}_{e}^{\mathcal{R}_{e}}(k)\right\rVert}{\bf p}_{\mathcal{R}_{ed}}^{\mathcal{R}_{e}}(k),$ (28)
where $K_{s}\in\mathbb{R}^{+}$ is a proportional gain representing a trade-off between stability and responsiveness of the scaling action. Indeed, $K_{s}$ close to 00 reduces sudden variations of $\underline{{\bf p}}_{\mathcal{R}_{ed}}^{\mathcal{R}_{e}}(k)$, even in presence of noisy force measures, but provides less force-limiting action. On the other hand, $K_{s}$ significantly greater than 00 makes the scaling action more responsive, but may yield oscillations in the presence of noisy measurements. Therefore, to increase the system’s robustness to noisy forces, exhibiting high-frequency variations, ${\bf f}_{e}^{\mathcal{R}_{e}}$ is filtered online as done for positions in (6). We note that the strategy in (28) does not allow forces to be bounded below a predefined limit. However, it prevents contact forces from growing indefinitely, ensuring a limitation even in the transients of the interaction and for any orientation of the end-effector, without relying on the knowledge of the dental environment’s complex geometry.
III-D Haptic feedback controller
In performing medical procedures, the operator relies heavily on the sense of touch. Therefore, in a HBTS providing realistic haptic sensations, it becomes critical to the success of the procedures. In impedance systems, the interaction is captured by both the contact force ${\bf f}_{e}^{\mathcal{R}_{e}}$ and the position error $\tilde{\bf p}^{\mathcal{R}_{e}}$, the latter being characterized by slower transients than the former, which exhibits a step dynamics. As already validated in a previous work [42], replacing ${\bf f}_{e}^{\mathcal{R}_{e}}$ with $\tilde{\bf p}^{\mathcal{R}_{e}}$ to generate the force feedback makes each contact state transparent to the operator, preventing them from having to deal with sudden force changes they are not used to from previous training and fieldworks. Therefore, the proposed HFC defines a second (virtual) impedance model, generating the force feedback from the motion error of the admittance controller:
${\bf f}_{h}^{\mathcal{H}_{b}}(k)={\bf K}_{h}\tilde{{\bf p}}^{\mathcal{H}_{b}}(k)+{\bf D}_{h}\dot{\tilde{{\bf p}}}^{\mathcal{H}_{b}}(k),$ (29)
where ${\bf K}_{h}$ and ${\bf D}_{h}$ respectively are virtual stiffness and damping diagonal matrices of the proper size. The position error $\tilde{{\bf p}}^{\mathcal{H}_{b}}(k)$ and its time derivative $\dot{\tilde{{\bf p}}}^{\mathcal{H}_{b}}(k)$ are defined, respectively, as:
$\tilde{{\bf p}}^{\mathcal{H}_{b}}(k)$ $={\bf R}_{\mathcal{H}_{e}}^{\mathcal{H}_{b}}(k)\left[{\bf R}_{\mathcal{H}_{e}}^{\mathcal{R}_{e}}(k)\right]^{\top}\tilde{{\bf p}}^{\mathcal{R}_{e}}(k),$ (30)
$\dot{\tilde{{\bf p}}}^{\mathcal{H}_{b}}(k)$ $={\bf R}_{\mathcal{H}_{e}}^{\mathcal{H}_{b}}(k)\left[{\bf R}_{\mathcal{H}_{e}}^{\mathcal{R}_{e}}(k)\right]^{\top}\dot{\tilde{{\bf p}}}^{\mathcal{R}_{e}}(k).$ (31)
We note that, as in (17), the rotation matrix ${\bf R}_{\bf z}^{\mathcal{H}_{e}}(\varphi(k))$ is used to provide the force feedback correspondence with the interaction observed in the camera view. With regard to the ${\bf K}_{h}$ and ${\bf D}_{h}$ parameters of the virtual system in (29), they are tuned to provide a natural force feedback to the human operator while the FR interacts with the environment, decoupling it from the admittance control dynamics. Specifically, the stiffness matrix ${\bf K}_{h}$ can be tuned to achieve the desired transient response, while the damping matrix ${\bf D}_{h}$ can be tuned to mitigate vibrations. We envisage the possibility that the generated force feedback may exceed the range of forces that the HD device can exert. Therefore, the HFC saturates the norm of the force feedback $\lVert{\bf f}_{h}^{\mathcal{H}_{b}}(k)\rVert$ to the maximum exertable force $f_{MAX}^{\mathcal{H}_{b}}\in\mathbb{R}^{+}$:
$\underline{\mathbf{f}}_{h}^{\mathcal{H}_{b}}(k)=\begin{cases}\mathbf{f}_{h}^{\mathcal{H}_{b}}(k),&\text{if}~{}\lVert\mathbf{f}_{h}^{\mathcal{H}_{b}}(k)\rVert<f_{MAX}^{\mathcal{H}_{b}};\\ f_{MAX}^{\mathcal{H}_{b}}\dfrac{\mathbf{f}_{h}^{\mathcal{H}_{b}}(k)}{\lVert\mathbf{f}_{h}^{\mathcal{H}_{b}}(k)\rVert},&\text{otherwise}.\end{cases}$ (32)
Finally, given the slow dynamics of $\tilde{{\bf p}}^{\mathcal{R}_{e}}$, the human operator may perceive residual force feedback even when the robot is no longer interacting. To address this, we introduce a force dead-band $f_{db}\in\mathbb{R}^{+}$ to prevent the HD from reacting to such residuals:
$\underline{{\bf f}}_{h}^{\mathcal{H}_{b}}(k)=\begin{cases}\underline{{\bf f}}_{h}^{\mathcal{H}_{b}}(k),&\text{if}~{}\left\lVert{\bf f}_{e}^{\mathcal{R}_{e}}(k)\right\rVert>f_{db};\\ {\bf 0},&\text{otherwise}.\end{cases}$ (33)
IV Experimental validation
To evaluate the performance of the proposed HBTS, we conduct interaction experiments using a phantom. Each experiment involves reaching a designated target tooth with an inactive dental scaler and applying a force to its surface, keeping the tool stationary. During the experiments, the robot remains out of the human operator’s view, who relies solely on the visual feedback from the camera and the force feedback from the HD. We select three target teeth: the lower right first premolar, the upper right central incisor, and the upper right first molar, as shown in Figure 6. The chosen targets are significant since each requires a different tool orientation, making the use of eye-hand coordination essential. For each target, we perform three interaction experiments with active force limitation and three without. With the aim of confirming the impact of the proposed strategy on the safety of interaction, we estimate, from the data recorded during the experiments, the conditional expectation of the interaction forces given a virtual penetration commanded by the operator. The conducted analysis is reported in Section IV-B.
Figure 6: Target teeth for dental scaling experiments: 1) lower right first premolar; 2) upper right central incisor; 3) upper right first molar.
In order to quantitatively assess the operation of the eye-hand coordination, we conduct a free-space experiment which concerns commanding the same spatial motion in the HD space, with different viewing angles. The results of the task execution are reported in Section IV-C.
Finally, we provide an experimental demonstration of the end-to-end system by conducting interaction experiments with the three target teeth in both scenarios, this time emulating the dental scaling procedure. The outcomes of a dental scaling procedure execution, representative of the overall results are shown in Section IV-D.
IV-A Experimental setup
The experimental setup involves the following equipment and materials:
  • a leader control station featuring a 6-degrees-of-freedom (DOF) force-feedback HD [50] and a workstation. The latter is a general purpose computer executing the leader controller loop, and communicating with the FR over an Ethernet network. In such a network, delays and packet losses are negligible, as is the case of an healthcare facility;
  • a FR consisting of the 6-DOF UR10 robotic arm by Universal Robots [51]. It is equipped with an integrated mechanical end-effector for free-hand dental procedures. Its design is discussed in Section IV-A1;
  • a dental phantom, i.e. 3D-printed life-sized model of a human mouth.
IV-A1 Integrated mechanical end-effector
In order to successfully perform any free-hand dental procedure, the operator relies on three basic factors: (a) the procedure-specific clinical tool; (b) the surgical field view, which may require external illumination and tools to show obstructed or poorly lit areas e.g., a mirror; (c) the sense of touch to ensure the right pressure to be exerted on the area of interest, in order to achieve a safe and effective procedure. In this work, we design an integrated mechanical end-effector, shown in Figure 7, that can provide all the tools and sensory capabilities required to perform free-hand dental procedures.
The designed end-effector consists of five main components:
  1. a)
    a 3D-printed mechanical interface connecting the robot’s flange with the force/torque sensor;
  2. b)
    a force/torque sensor measuring contact forces. The sensor used is the FT150 6-axis force/torque sensor by Robotiq [52];
  3. c)
    a 3D-printed tool socket supporting the clinical tool and the endoscopic camera. The interface consists of a rigid structure to enable proper transmission of forces from the tool to the sensor, a circular hole, adjustable via a bolt-and-nut mechanism, to accommodate tools of different sizes, and a movable camera-support ring, allowing for adjustments of the view and light beam angle;
  4. d)
    an HD endoscopic camera, equipped with LED illumination to ensure good visibility even in poorly lit areas;
  5. e)
    a clinical tool for performing the procedure of interest. The tool is placed in the supporting socket after being wrapped with friction tape to ensure good mechanical compliance, emulating the skin friction between the operator’s hand and the tool. The tool used for the chosen use case is a scaler.
Figure 7: Integrated mechanical end-effector for free-hand procedures, consisting of: 1) 3D-printed mechanical interface connecting the robot’s flange with the force/torque sensor; 2) force/torque sensor; 3) 3D-printed tool socket supporting the clinical tool and the endoscopic camera; 4) HD endoscopic camera; 5) clinical tool. Visible in the left side of the figure is also the dental phantom used.
IV-A2 Eye-hand coordination controller parametrization
[…]
With regard to the scaling factors, we set $\beta=1$ and ${\boldsymbol{\Lambda}}=diag(0.8,0.8,0.8,1.0,1.0,1.0)$ to achieve high control accuracy on the translational axes without significantly reducing the workspace size.
IV-A3 Admittance controller parametrization
Since the interaction experiments require modest compliance and high damping along the normal axis, and greater and more precise positioning along the tangential axes, we set the impedance parameters reported in Table I.
Axis of ${\mathcal{R}_{e}}$ ${\bf K}_{P}(N/m)$ ${\bf K}_{D}(Ns/m)$ ${\bf M}_{d}(kg)$
x $50000$ $2500$ $50$
y $50000$ $2500$ $50$
z $6000$ $12000$ $50$
TABLE I: Admittance controller parameters
Regarding the force limitation strategy, we identify a suitable trade-off between responsiveness and stability by setting $K_{s}=0.06$, and by filtering the exerted forces with a moving average filter of window size $n=500$.
IV-A4 Virtual feedback parametrization
With regard to the virtualization of the force feedback, the virtual impedance matrices are tuned to provide the human operator with a natural sensation, ensuring both smooth transient response and reduced vibrations. Therefore, we set the parameters reported in Table II.
[…]
Axis of ${\mathcal{H}_{b}}$ ${\bf K}_{h}(N/m)$ ${\bf D}_{h}(Ns/m)$
x $0.1$ $0.5$
y $0.1$ $0.5$
z $0.1$ $0.5$
TABLE II: Virtual feedback parameters.
IV-B Force limitation assessment
[…]
Given the conditional probability function $p\left(a(k)|b(k)\right)$, we estimate the conditional expectation $E\left(A|B=b_{d}\right),\forall b_{d}\in\mathcal{B}_{d}$:
$E\left(A|B=b_{d}\right)=\sum_{a_{d}\in\mathcal{A}_{d}}a_{d}~{}p(A=a_{d}|B=b_{d}),$ (34)
where $a_{d}$ is the discrete value of a bin in $\mathcal{A}_{d}$, and the conditional variance $\text{Var}(A|B=b_{d}),\forall b_{d}\in\mathcal{B}_{d}$:
$\begin{split}\text{Var}(A|B=b_{d})=&\sum_{a_{d}\in\mathcal{A}_{d}}(a_{d}-E(A|B=b_{d}))^{2}\\ &\quad\cdot p(A=a_{d}|B=b_{d}).\end{split}$ (35)
In Figure 8 are reported the conditional expectations of the exerted force $E\left(f_{e}^{\mathcal{R}_{e},\perp}\middle|\left\lVert{\bf p}_{\mathcal{R}_{ed}}^{\mathcal{R}_{e}}\right\rVert\right)$ and the conditional standard deviation, obtained as the square root of the conditional variance, in both scenarios.
Figure 8: Comparison between the conditional expectation of the interaction force in Scenario A and Scenario B.
To evaluate the effectiveness of the force limitation strategy, we compared $E\left(A|B=b_{d}\right),\forall b_{d}\in\mathcal{B}_{d}$ and $\text{Var}(A|B=b_{d}),\forall b_{d}\in\mathcal{B}_{d}$ in both scenarios. To test whether the conditional variances differ significantly, we perform Levene’s test and Bartlett’s test. Levene’s test shows that the conditional variances are significantly different (p-value $p\approx 0$), indicating that force limitation strategy affects the dispersion of normal force $f_{e}^{\mathcal{R}_{e},\perp}(k)$ with respect to virtual penetration $\left\lVert{\bf p}_{\mathcal{R}_{ed}}^{\mathcal{R}_{e}}(k)\right\rVert$. Bartlett’s test, which is more sensitive to normal distributions, confirms this observation ($p\approx 0$). This suggests that the difference in variance is not only due to the sampling effect but is a structural feature of the collected data. We compare the conditional expectation using a t-test for independent samples. Since Levene’s test suggests a significant variance difference, we apply Welch’s t-test, which is more robust in the case of unequal variances. The t-test returned a t-statistic value of $t=-6.18$, confirming that the normal force is lower in Scenario A, and $p\approx 0$. Since such value is below the significance threshold $\alpha=0.05$, we can reject the null hypothesis and conclude that the conditional expectations of normal force in the two scenarios are significantly different.
IV-C Eye-hand coordination assessment
In order to perform a formal assessment of eye-hand coordination, we execute a teleoperated task in the free space. The human operator performs a translatory motion with the HD, simultaneously rotating the viewing angle by rotating the HD’s stylus by an angle $\varphi$ about the $\bf z$ axis of $\mathcal{H}_{e}$. The validation is considered completed if the motion transferred to the robot varies as the viewing angle does. Figure 9 shows the $\varphi$ angle, the position of HD’s stylus and the position of the FR, respectively.
(a) $\varphi$ angle about the $z$ axis of $\mathcal{H}_{e}$.
(b) Position of HD’s stylus.
(c) Position of the FR.
Figure 9: Eye-hand coordination assessment.
[…]
These results confirm that the mapping of motion between HD space and FR space is affected by the rotation of the visual field. However, they do not provide any suggestion regarding the effectiveness of eye-hand coordination for the human operator. Therefore, in we provide a comprehensive qualitative assessment of the latter.
IV-D End-to-end system experimental demonstration
We provide a demonstration of the end-to-end system’s operation in the accompanying video , where we showcase the execution of the dental scaling procedure on the three target teeth in Scenario A. In this section, we compare the outcomes of the execution of a dental scaling procedure in Scenario A, against those obtained in Scenario B. To avoid repetition, the analysis focuses on the outcomes of executing a dental scaling action on a single target tooth, the lower right first premolar.
In Figure 10 the normal forces exerted by the FR in both scenarios are shown. A visual inspection showcases a significant reduction in interaction forces when the free-hand force limitation strategy is active, with a decrease of approximately $50\%$ in Scenario A, thereby confirming the validation presented in Section IV-B. The effectiveness of this strategy is made even more evident by Figure 11, which shows the norm of the references provided to the admittance controller. Notably, in Scenario A, where the interaction forces are lower, the virtual penetration commanded to the controller is even greater than in Scenario B. Nevertheless, as shown in Figure 11a, the implementation of this strategy enables the scaling of the provided references, effectively reducing the contact forces and enhancing the system’s safety.
(a) Contact forces in Scenario A.
(b) Contact forces in Scenario B.
Figure 10: Nomal forces in both scenarios.
(a) Norm of the admittance controller references in Scenario A.
(b) Norm of the admittance controller references in Scenario B.
Figure 11: Norm of the admittance controller references in both scenarios.
Finally, in Figure 12 the forces rendered on the HD in both scenarios are shown. As expected, the results are consistent between the two scenarios, while both showing high similarity to exerted force profiles as well as slow dynamics. Such aspects enhance naturalness for the human operator.
(a) Forces rendered on the HD in Scenario A.
(b) Forces rendered on the HD in Scenario B.
Figure 12: Forces rendered on the HD in both scenarios.
V Discussions and conclusions
Motivated by a lack of research on robotized free-hand dental procedures, this paper proposed a HBTS specifically designed to this aim. The system features visual and haptic feedback, that are appropriately combined and coordinated to allow for a natural teleoperation experience. To improve the visibility of the intervention site, we designed a dedicated end-effector accommodating an eye-in-hand endoscopic camera, and ensuring compatibility with existing clinical tools. We improved the accuracy of the procedure by introducing human tremor filtering and dynamic scaling of velocities. By providing the eye-hand coordination and the TCP control, we enabled the human operator to perceive the motion correspondence between their movements and those of the end-effector, observed in the camera view, preserving manual dexterity and expertise. Additionally, by appropriately transforming the virtualized force feedback, we also ensured the force correspondence, enabling the human operator to experience a highly natural sensory perception. Finally, we enabled safe interactions, by scaling the motion references provided to the admittance controller. The proposed strategy allows for accurate force limitation in any contact state and for any orientation of the interaction tool, without requiring any prior modeling of the intervention site.
We believe that our HBTS is a necessary intermediate step for a broader adoption of robotics in dentistry, as it allows for accurate treatments, while preserving the operator’s cognitive and technical abilities. With respect to manual procedures, our system allows to preserve operator’s ergonomics, track operator’s performance, and enable remote healthcare.
We emphasize that the proposed system allows for the collection of haptic and motion data, which can be crucial during the training phase of new operators for the evaluation of learning metrics, i.e learning curves [55], or for the derivation of data-driven control policies [56] for autonomous robots.
With our future work we intend to extend the system with autonomous patient motion compensation to further enhance the system’s safety. Likewise, we intend to conduct experiments with professionals and active dental tools both to confirm the system’s usability under more representative clinical conditions and to collect expert policies for (semi-)autonomous procedures.
Acknowledgments
The authors would like to thank Sabatino Panico for contributing to the system calibration and experimental campaigns.

### v2

@@
 Free-hand dental procedures are typically repetitive, time-consuming and require high precision and manual dexterity.
-Dental robots can play a key role in improving procedural accuracy and safety, enhancing patient comfort, and reducing operator workload.
+Robots can play a key role in improving procedural accuracy and safety, enhancing patient comfort, and reducing operator workload.
 However, robotic solutions for free-hand procedures remain limited or completely lacking, and their acceptance is still low.
-To address this gap, we develop a haptic bilateral teleoperation system (HBTS) for free-hand dental procedures.
+To address this gap, we develop a haptic bilateral teleoperation system (HBTS) for free-hand dental procedures (FH-HBTS).
 The system includes a dedicated mechanical end-effector, compatible with standard clinical tools, and equipped with an endoscopic camera for improved visibility of the intervention site.
@@
 This ensures effective force limitation in all contact states without requiring prior knowledge of the environment.
-The proposed HBTS is validated in a dental scaling procedure using a dental phantom.
+The proposed FH-HBTS is validated in a dental scaling procedure using a dental phantom.
 The results show that the system improves the naturalness, safety, and accuracy of teleoperation, highlighting its potential to enhance free-hand dental procedures.
@@
 I Introduction
-I-A Background
-Robotics is rewriting the future of healthcare, emerging as a disruptive technology capable of optimizing and revolutionizing the way some medical procedures are conducted.
-Technological advancements enable robots to play an increasingly prominent role in the execution of complex medical procedures, where the required scale, accuracy, and precision are often beyond human capabilities.
-The successful application of such technology in the medical field is feeding a wave of innovation in dentistry, where robotics is changing the game [1].
-Figure 1: Manual vs teleoperated dental procedures (pictorial views). In the upper right corner, is the console of The Senhance® Surgical System [2].
 The World Health Organization estimates that oral diseases affect nearly 3.5 billion people worldwide.
 However, access to primary oral health services is limited due to the high costs of procedures, the unequal distribution of oral health professionals, and the lack of adequate healthcare facilities [1].
-Furthermore, many procedures are conducted free-hand by uninformed operators, resulting in highly inefficient and disastrous outcomes.
 Traditionally, routine, prosthetic, periodontal, and some surgical dental procedures are performed free-hand.
@@
 Periodontal procedures include scaling and root planing (SRP), which involve the removal of plaque and calculus from the surfaces of the teeth and the smoothing of root surfaces, typically employing manual hand instruments and inserts mounted on sonic or ultrasonic handpieces [4].
-Finally, although computer-assisted implant surgery (s-CAIS), dynamic computer-assisted implant surgery (d-CAIS), and robotic computer-assisted implant surgery (r-CAIS) have shown a significant increase in the accuracy of dental implant placement [7], many operators still prefer the free-hand method for dental implantology [8].
-Each of these procedures, even the simplest one, requires accuracy, manual skills and dexterity on the part of the operator.
-Indeed, the latter has to cope with a highly unstructured environment, which is full of natural obstacles that occlude the view of the surgical site and reduce the space for maneuvering, requiring constant changes in the orientation of the surgical tool and the operator’s posture.
-The successful outcome of a procedure, as well as the comfort for the patient, therefore, depend on both human factors, i.e., the experience and skills of the human operator [9], the adequacy of the instrumentation used and the duration of the treatments [10], and environmental factors, i.e., difficult anatomical conditions and less accessibility to the surgical site [11].
+Finally, although robotic computer-assisted implant surgery (r-CAIS) have shown a significant increase in the accuracy of dental implant placement [5], many operators still prefer the free-hand method for dental implantology [6].
+Figure 1: Manual (left) vs teleoperated (right) dental procedures (pictorial views).
+Each of these procedures, even the simplest ones, require accuracy, extensively trained manual skills and dexterity.
+Remark 1.
+The oral cavity is a highly unstructured environment, which is full of natural obstacles that occlude the view of the surgical site and reduce the space for maneuvering, requiring constant changes in the orientation of the clinical tool and the operator’s posture.
+Therefore, the successful outcome of a procedure, as well as the comfort for the patient, depend on both human factors, i.e., the experience and skills of the human operator [7], the adequacy of the instrumentation used and the duration of the treatments [8], and environmental factors [9], as highlighted in Remark 1.
 In particular, the latter factors can significantly affect the operator’s comfort.
@@
 The main causes of these disorders can be traced back to an incorrect application of ergonomics, such as awkward postures, repetitive and prolonged movements, use of small or narrow tool handles, exertion of forces on hard or sharp objects, and exposure to vibrations [11].
-Therefore, in the process of training new operators, much emphasis is placed on learning skills related to space utilization and positioning of surgical and non-surgical instruments, e.g., mirrors, to overcome occlusions, while maintaining healthy and comfortable posture [14].
+Hence, when training new operators, much emphasis is placed on learning skills related to space utilization and positioning of surgical and non-surgical instruments, e.g., mirrors, to overcome occlusions, while maintaining healthy and comfortable posture [12].
 In addition to this, to ensure the safety and effectiveness of the procedure, new operators must acquire skills related to the correct amount of pressure to apply to the site of interest, depending on the type of tool used.
-Consequently, the sense of touch plays a key role in achieving successful procedures.
-This is made evident by emerging trends in teaching, which make use of virtual environments and haptic devices (HD)s for training new operators [15, 16].
-In contrast to humans, robots are capable of achieving accuracy, precision, efficiency and scales beyond human capabilities.
-For this reason, the role of dental robots is to improve dental procedures, both in terms of patient satisfaction, i.e., broader access to care, reduced treatment time, successful outcome of the procedure, fast recovery, and in terms of operator comfort, i.e., reduced workload intensity, increased visibility of the surgical field, improved ergonomics [17], as shown in Figure 1.
-Haptic teleoperations are a technology that combines the robot’s skills with the flexibility, dexterity, and cognitive capabilities of humans.
-This technology is applied in multiple domains, such as telesurgery and space exploration, as it overcomes barriers in terms of distance and hazardous environments, and scales beyond human reach.
+Remark 2.
+The sense of touch plays a key role in achieving successful and efficient procedures.
+This is also apparent from emerging trends in teaching, which make use of virtual environments and haptic devices (HD)s for training new operators [13, 14].
+Robots can achieve accuracy, precision, efficiency beyond human capabilities.
+For this reason, the role of robots is to improve dental procedures, both in terms of patient satisfaction, i.e.
+broader access to care, reduced treatment time, fast recovery, and in terms of operator comfort, i.e., reduced workload intensity, increased visibility of the surgical field, improved ergonomics [15], as shown in Figure 1.
+I-A Related works
+Haptic teleoperations are a technology that enables the remote execution of complex tasks, employing a robotic system.
+Remark 3.
+Haptic teleoperations combine the robot skills with the flexibility, dexterity, and cognitive capabilities of humans, while overcoming barriers in terms of distance, hazardous environments, and scales beyond human reach.
 By providing multiple types of sensory information, i.e.
-visual and haptic feedback, these systems enable the human operator to experience telepresence, defined as the perception of interacting with the remote environment as if it were directly encountered [18].
-Furthermore, by integrating different haptic feedback and guidance modalities [19, 20], and shared control approaches [21], they contribute to reducing errors and enhance the efficiency, accuracy and safety of procedures, thus extending human capabilities, while also improving operator comfort and confidence.
-I-B Related works
+visual and haptic feedback, HBTSs enable the human operator to experience telepresence, i.e.
+the perception of interacting with the remote environment as if it were directly encountered [16].
+Furthermore, by integrating different haptic feedback and guidance modalities [17, 18], and shared control approaches [19], they contribute to reducing errors and enhance the efficiency, accuracy and safety of procedures, while also improving operator comfort and confidence.
+For these reasons, such technology is applied in multiple domains, such as telesurgery, space exploration, and the handling of hazardous materials in nuclear or underwater environments.
+However, in the context of dentistry, the adoption of such systems is still very limited.
 The majority of research in robotics for dentistry is focused on oral implantology [20].
-Robotic oral implant surgery offers significant benefits in terms of implant accuracy in partially [23] or totally endetulous patients, because of the control of angular deviation and axial errors, also compared to a s-CAIS system [24].
-Better placement accuracy, however, can be achieved at the expense of surgical efficiency, as shown in [25], which compares surgical efficiency and implant positioning accuracy among three types of robotic systems: passive, semi-active, and active robots.
-Such technologies also enable flapless dental implant placement, speeding up recovery time, reducing postoperative pain and swelling, and enhancing patient comfort.
+Robotic oral implant surgery offers significant benefits in terms of implant accuracy in partially [21] or totally edentulous patients, also compared to computer-assisted implant surgery (s-CAIS) systems [22].
+Better placement accuracy, however, can be achieved at the expense of surgical efficiency, as shown in [23].
+Such technologies also enable flapless dental implant placement, speeding up recovery time, reducing postoperative pain, and enhancing patient comfort [24].
 [25] shows how a telerobotic-assisted drilling system can enhance the accuracy of the human operator while preserving awareness of the procedure through an effective transmission of cutting forces and vibrations.
-On the other hand, the state-of-the-art shows considerably less research interest in other free-hand procedures [1].
+On the other hand, the state-of-the-art shows considerably less research effort in other free-hand procedures [26].
 In the context of prosthetic procedures, robotics can enhance the precision of manual grinding using a turbine drill during tooth preparation [27] and can also compensate for hand tremors and the lack of tactile feedback in laser ablation [28, 29], improving accuracy [30] and efficiency [31].
 In endodontics, robotics can automate and enhance precision and efficiency in root canal treatments [32, 33], outperforming current approaches also in effectively addressing biological and mechanical traits associated with biofilm recalcitrance and reinfection [34].
-In periodontal procedures, robotic systems have the potential to increase accuracy, reducing the risk of errors and trauma to tissues by optimizing applied force, as with other free-hand procedures.
+In periodontal procedures, robotic systems have the potential to increase accuracy, reducing the risk of errors and trauma to tissues by optimizing applied forces, as with other free-hand procedures.
 They can also improve the ergonomics for the operator, and comfort for the patient by speeding up recovery time.
-However, for such procedures robotic technologies are still lacking [35], and there is no evidence of such advantages.
-Thus, although robotics emerges as a transformative technology in the context of dental practice, there are still many limitations to consider.
-First of all, the quality of research is still insufficient in terms of clinical validation and technology readiness level (TRL) [36].
+However, for such procedures, robotic technologies are still lacking [35], and there is no empirical evidence of such advantages.
+Thus, although robotics shows promise in dental practice, several limitations remain.
+First, the quality of research is still insufficient in terms of clinical validation and technology readiness level (TRL) [36].
 This is heavily influenced by the high costs of these robotic systems and the need to ensure their compatibility with existing clinical tools [37].
-Second, the most widely used category of robots is fully autonomous ones.
-However, the use of this type of robots requires the development of state-of-the-art algorithms and advanced capabilities in perception, recognition, and exception handling [37].
-In this context, the presence of a human, who must possess digital and mechanical skills in addition to clinical expertise, is still necessary not only as a supervisor but also as an operator.
-Considering the absence of haptic feedback in most robotic systems, which would improve the efficiency and effectiveness of procedures and increase operator acceptance [38], as well as the need to provide access to care in areas with a shortage of professionals, haptic teleoperations represent a viable solution.
-These systems would enable the remote execution of procedures with the accuracy and efficiency of robots while preserving human cognitive abilities.
-Moreover, they would allow the acquisition of human control strategies directly exercised in the robot state space, simplifying the process of transferring the knowledge from the human to the robot [39] and enabling the effective and objective evaluation of operator skills during training [40].
-Finally, research tends to prioritize oral implantology over other dental procedures, despite the high number of free-hand routine procedures performed daily and patient preferences, as they tend to accept these technologies more readily when the degree of invasiveness is lower [41].
-Considering these limitations, [1] recommends to shift the focus from oral implantology procedures towards less invasive ones, which are better suited as forerunners in the field, and also to invest in educational robotics research in academia as a valuable tool to introduce dental robotics and overcome the barrier of acceptance of robotic systems among future dentists.
-I-C Contributions
-In our previous work, we proposed an HBTS to facilitate the execution of precision tasks and increase the safety of interaction using haptic teleoperations [42].
-Motivated by the limitations or complete absence of robotic solutions for the most repetitive free-hand dental procedures, and recognizing the importance of haptic teleoperations in enhancing human capabilities while preserving dexterity, cognitive skills, and procedural awareness, in this paper we present an extended version of the system.
-Our contribution is an HBTS for free-hand dental procedures, including:
+Second, the most widely used robots are fully autonomous, which require the development of state-of-the-art algorithms and advanced capabilities in perception, recognition, and exception handling [37].
+In this context, the presence of a human is still necessary not only as a supervisor but also as an operator.
+In light of Remark 2 and Remark 3, and considering the impact on the operator acceptance [38], haptic teleoperations represent a viable solution.
+Moreover, the HBTSs would allow the acquisition of human control strategies directly exercised in the robot state space, simplifying the process of transferring the knowledge from the human to the robot [39] and enabling the effective and objective evaluation of operator skills during training [40].
+Finally, research tends to prioritize oral implantology over other dental procedures, despite the high number of daily free-hand routine procedures and the less invasiveness, entailing better patient acceptance [41].
+Considering these limitations, [26] recommends to shift the focus from oral implantology procedures towards less invasive ones, which are better suited as forerunners in the field, and also to invest in educational robotics research in academia as a valuable tool to overcome the barrier of acceptance of robotic systems among future dentists.
+I-B Contributions
+Motivated by the limitations or complete absence of robotic solutions for the most repetitive free-hand dental procedures, and in light of Remark 3, in this paper we present a HBTS for free-hand dental procedures (FH-HBTS), including:
   • a dedicated mechanical end-effector, compatible with existing clinical tools and featuring an endoscopic camera, to improve visibility at the intervention site;
   • a virtualized force feedback, consistent with the interaction observed through the visual feedback of the eye-in-hand camera, which enhances the haptic experience for the operator;
-  • a control system that ensures procedural safety, improves accuracy, and preserves manual dexterity, while providing motion correspondence between the operator’s movements and those of the end-effector, as observed in the camera’s view.
-We validate the proposed HBTS on a free-hand procedure, consisting of dental scaling.
+  • a control system that ensures procedural safety, improves accuracy, and preserves manual dexterity, while providing motion correspondence between the operator’s movements and those of the end-effector, as observed in the camera view.
+To the best of the authors’ knowledge, this is the only HBTS designed for free-hand dental procedures. Key contributions are:
+  • a control system to ensure accurate control of the tool center point (TCP) and to guarantee motion correspondence, using a novel eye-hand coordination strategy;
+  • a force limitation strategy that makes no previous assumption or estimation of the environment geometry, therefore accounting for the challenges outlined in Remark 1;
+  • eye-hand force correspondence provided with saturation and filtering via a deadband function;
+  • TRL-4 validation of the FH-HBTS on a free-hand procedure, consisting of dental scaling.
 II Preliminaries
@@
   • $\mathcal{R}_{b}$ is the base reference frame, in which motion references are expressed;
-  • $\mathcal{R}_{e}$ is the robot’s end-effector frame (computed from forward kinematics of joint measurements), corresponding to the tool center point (TCP);
+  • $\mathcal{R}_{e}$ is the robot’s end-effector frame (computed from forward kinematics of joint measurements), corresponding to the TCP;
   • $\mathcal{R}_{ed}$ is the desired robot’s end-effector frame, input to admittance control;
@@
 III Methodology
-III-A System overview
-The HBTS we propose, shown in Figure 3, is based on a position-force architecture [43], coupling the leader (LR) and follower (FR) robots.
+The FH-HBTS we propose, shown in Figure 3, is based on a position-force architecture [42], coupling the leader (LR) and follower (FR) robots.
 The LR is an impedance HD, sensing the pose of the human operator and rendering haptic feedback according to the remote interaction between the FR and the environment.
@@
 Both controllers record data throughout the entire procedure.
-A detailed description of the objectives and operation of the two controllers is provided in Section III-B and Section III-D, respectively.
 The FR is a position-controlled robot, instructed to track the HD through an admittance controller, which provides the latter with a configurable degree of compliance on its target pose with respect to the environment.
@@
 Finally, it limits the interaction forces by scaling the received motion references through a free-hand force limitation strategy embedded in the controller itself.
-The force limitation strategy is explained in Section III-C2.
+With reference to Figure 3, depicting the complete system, the following sections examine its individual subcomponents in detail.
+In Section III-A and Section III-C, a detailed description of the objectives and operation of the EHCC and HFC controllers is provided.
+In Section III-B the robot interaction control is described.
 Figure 3: Haptic bilateral teleoperation system scheme.
-III-B Eye-hand coordination controller
-In an HBTS, visual feedback plays a crucial role in overcoming the limitations caused by distance.
+III-A Eye-hand coordination controller
+In a HBTS, visual feedback plays a crucial role in overcoming the limitations caused by distance.
 In teleoperated free-hand dental procedures, using an eye-in-hand camera is the most effective solution [43].
@@
 This creates a mismatch between the movements expected by the operator and those actually executed by the robot in the camera view, compromising the accuracy of the task.
-The da Vinci Surgical SystemTM, which requires the insertion of multiple tools to implement an eye-to-hand strategy, provides the motion correspondence between the operator’s movements and those of the end-effector, observed in the camera view, by projecting the endoscope image above the surgeon’s hands by means of mirrored overlay optics [45].
+The da Vinci Surgical SystemTM, which requires the insertion of multiple tools to implement an eye-to-hand strategy, thereby increasing the bulk of the equipment and the overall costs, provides the motion correspondence between the operator’s movements and those of the end-effector, observed in the camera view, by projecting the endoscope image above the surgeon’s hands by means of mirrored overlay optics [44].
 In our eye-hand coordination controller, we provide the motion correspondence at the command level, by creating a coupling between the HD’s stylus and the robot’s TCP, and appropriately converting the motion of the former into the motion of the FR.
@@
 This includes defining the desired relative rotation ${\bf R}_{\mathcal{H}_{e}}^{\mathcal{R}_{e}}$, and performing manually the alignment.
-This is done by the human operator during what we indicate as engagement phase.
-III-B1 Engagement phase
+This is done by the human operator during what we indicate as the engagement phase.
+The latter is presented in Section III-A1, while the human tremors filtering and the eye-hand coordination mapping, implementing the motion correspondence, are discussed in Section III-A2 and Section III-A3, respectively.
+III-A1 Engagement phase
 During the engagement phase, the EHCC retrieves the homogeneous transformation matrix of the HD’s stylus ${\bf T}_{\mathcal{H}_{e}}^{\mathcal{H}_{b}}(k)$, and converts it into a robot’s end-effector homogeneous transformation matrix with the following kinematic equations:
@@
 Figure 4: Engagement phase supported by the visualization tool.
-III-B2 Human tremors filtering
+III-A2 Human tremors filtering
 At each control cycle, the EHCC retrieves the actual pose ${\bf x}_{\mathcal{H}_{e}}^{\mathcal{H}_{b}}(k)$ of the HD’s stylus, corresponding to ${\bf T}_{\mathcal{H}_{e}}^{\mathcal{H}_{b}}(k)$.
@@
 Keeping the cumulative sums ${\bf s}(k)$ and ${\bf c}(k)$ enables achieving $\mathcal{O}(1)$ computational complexity.
-III-B3 Eye-hand coordination
-In an HBTS, coupling two robots with different kinematics requires the use of mapping and scaling techniques.
+III-A3 Eye-hand coordination
+In a HBTS, coupling two robots with different kinematics requires the use of mapping and scaling techniques.
 The literature shows that this coupling can be achieved by converting the motion of the HD into that of the FR through simple kinematic transformations, and appropriately scaling it to either expand the workspace size of the HD or enhance control accuracy [46].
@@
 6:Extract ${\bf x}_{\mathcal{R}_{ed}}^{\mathcal{R}_{b}}(k)$ from ${\bf T}_{\mathcal{R}_{ed}}^{\mathcal{R}_{b}}(k)$
-III-C Robot interaction control
-III-C1 Admittance model
+III-B Robot interaction control
+In this section, we first formalize the contact model between the robot and the environment (Section III-B1), and subsequently, we present the free-hand strategy for limiting the interaction forces (Section III-B2).
+III-B1 Admittance model
 Let ${\bf x}_{\mathcal{R}_{ed}}^{\mathcal{R}_{b}}(k)=\begin{bmatrix}{\bf p}_{\mathcal{R}_{ed}}^{\mathcal{R}_{b}}(k)^{\top}&{\bf q}_{\mathcal{R}_{ed}}^{\mathcal{R}_{b}}(k)^{\top}\end{bmatrix}^{\top}$ be the desired pose and $\mathcal{R}_{c}$ the compliant frame specified by the quantities ${\bf p}_{\mathcal{R}_{c}}^{\mathcal{R}_{e}}(k)$, $\dot{\bf p}_{\mathcal{R}_{c}}^{\mathcal{R}_{e}}(k)$, $\ddot{\bf p}_{\mathcal{R}_{c}}^{\mathcal{R}_{e}}(k)$ and ${\bf q}_{\mathcal{R}_{c}}^{\mathcal{R}_{e}}(k)$.
@@
 ${\bf p}_{\mathcal{R}_{c}}^{\mathcal{R}_{b}}(k)={\bf p}_{\mathcal{R}_{e}}^{\mathcal{R}_{b}}(k)+{\bf R}_{\mathcal{R}_{e}}^{\mathcal{R}_{b}}(k){\bf p}_{\mathcal{R}_{c}}^{\mathcal{R}_{e}}(k).$ (25)
-III-C2 Free-hand force limitation strategy
+III-B2 Free-hand force limitation strategy
 Assuming that the environment can be approximated by a linear spring model, the exerted force at the steady state [47] is:
@@
 However, it prevents contact forces from growing indefinitely, ensuring a limitation even in the transients of the interaction and for any orientation of the end-effector, without relying on the knowledge of the dental environment’s complex geometry.
-III-D Haptic feedback controller
+III-C Haptic feedback controller
 In performing medical procedures, the operator relies heavily on the sense of touch.
@@
 In impedance systems, the interaction is captured by both the contact force ${\bf f}_{e}^{\mathcal{R}_{e}}$ and the position error $\tilde{\bf p}^{\mathcal{R}_{e}}$, the latter being characterized by slower transients than the former, which exhibits a step dynamics.
-As already validated in a previous work [42], replacing ${\bf f}_{e}^{\mathcal{R}_{e}}$ with $\tilde{\bf p}^{\mathcal{R}_{e}}$ to generate the force feedback makes each contact state transparent to the operator, preventing them from having to deal with sudden force changes they are not used to from previous training and fieldworks.
+As proposed in [49], replacing ${\bf f}_{e}^{\mathcal{R}_{e}}$ with $\tilde{\bf p}^{\mathcal{R}_{e}}$ to generate the force feedback makes each contact state transparent to the operator, preventing them from having to deal with sudden force changes they are not used to from previous training and fieldworks.
 Therefore, the proposed HFC defines a second (virtual) impedance model, generating the force feedback from the motion error of the admittance controller:
@@
 IV Experimental validation
-To evaluate the performance of the proposed HBTS, we conduct interaction experiments using a phantom.
+To evaluate the performance of the proposed FH-HBTS, we conduct interaction experiments using a phantom.
 Each experiment involves reaching a designated target tooth with an inactive dental scaler and applying a force to its surface, keeping the tool stationary.
 During the experiments, the robot remains out of the human operator’s view, who relies solely on the visual feedback from the camera and the force feedback from the HD.
-We select three target teeth: the lower right first premolar, the upper right central incisor, and the upper right first molar, as shown in Figure 6.
+We select three target teeth: the lower right first premolar, the upper right central incisor, and the upper right second premolar, as shown in Figure 6.
 The chosen targets are significant since each requires a different tool orientation, making the use of eye-hand coordination essential.
@@
 The conducted analysis is reported in Section IV-B.
-Figure 6: Target teeth for dental scaling experiments: 1) lower right first premolar; 2) upper right central incisor; 3) upper right first molar.
+Figure 6: Target teeth for dental scaling experiments: 1) lower right first premolar; 2) upper right central incisor; 3) upper right second premolar.
 In order to quantitatively assess the operation of the eye-hand coordination, we conduct a free-space experiment which concerns commanding the same spatial motion in the HD space, with different viewing angles.
@@
 IV-A3 Admittance controller parametrization
-Since the interaction experiments require modest compliance and high damping along the normal axis, and greater and more precise positioning along the tangential axes, we set the impedance parameters reported in Table I.
-Axis of ${\mathcal{R}_{e}}$ ${\bf K}_{P}(N/m)$ ${\bf K}_{D}(Ns/m)$ ${\bf M}_{d}(kg)$
-x $50000$ $2500$ $50$
-y $50000$ $2500$ $50$
-z $6000$ $12000$ $50$
-TABLE I: Admittance controller parameters
+Since the interaction experiments require modest compliance and high damping along the normal axis, and greater and more precise positioning along the tangential axes, we set ${\bf K}_{P}=\text{diag}(50000,50000,6000)$, ${\bf K}_{D}=\text{diag}(2500,2500,12000)$ and ${\bf M}_{d}=\text{diag}(50,50,50)$.
 Regarding the force limitation strategy, we identify a suitable trade-off between responsiveness and stability by setting $K_{s}=0.06$, and by filtering the exerted forces with a moving average filter of window size $n=500$.
@@
 With regard to the virtualization of the force feedback, the virtual impedance matrices are tuned to provide the human operator with a natural sensation, ensuring both smooth transient response and reduced vibrations.
-Therefore, we set the parameters reported in Table II.
+Therefore, we set ${\bf K}_{h}=\text{diag}(0.1,0.1,0.1)$, ${\bf D}_{h}=\text{diag}(0.5,0.5,0.5)$.
 […]
-Axis of ${\mathcal{H}_{b}}$ ${\bf K}_{h}(N/m)$ ${\bf D}_{h}(Ns/m)$
-x $0.1$ $0.5$
-y $0.1$ $0.5$
-z $0.1$ $0.5$
-TABLE II: Virtual feedback parameters.
 IV-B Force limitation assessment
 […]
-Given the conditional probability function $p\left(a(k)|b(k)\right)$, we estimate the conditional expectation $E\left(A|B=b_{d}\right),\forall b_{d}\in\mathcal{B}_{d}$:
-$E\left(A|B=b_{d}\right)=\sum_{a_{d}\in\mathcal{A}_{d}}a_{d}~{}p(A=a_{d}|B=b_{d}),$ (34)
-where $a_{d}$ is the discrete value of a bin in $\mathcal{A}_{d}$, and the conditional variance $\text{Var}(A|B=b_{d}),\forall b_{d}\in\mathcal{B}_{d}$:
-$\begin{split}\text{Var}(A|B=b_{d})=&\sum_{a_{d}\in\mathcal{A}_{d}}(a_{d}-E(A|B=b_{d}))^{2}\\ &\quad\cdot p(A=a_{d}|B=b_{d}).\end{split}$ (35)
-In Figure 8 are reported the conditional expectations of the exerted force $E\left(f_{e}^{\mathcal{R}_{e},\perp}\middle|\left\lVert{\bf p}_{\mathcal{R}_{ed}}^{\mathcal{R}_{e}}\right\rVert\right)$ and the conditional standard deviation, obtained as the square root of the conditional variance, in both scenarios.
+Given the conditional probability function $p\left(a(k)|b(k)\right)$, we estimate the conditional expectation $E\left(A|B\right)$, and the conditional variance $\text{Var}(A|B)$ following [55].
+In Figure 8 are reported the conditional expectations of the exerted force $E\left(f_{e}^{\mathcal{R}_{e},\perp}\middle|\left\lVert{\bf p}_{\mathcal{R}_{ed}}^{\mathcal{R}_{e}}\right\rVert\right)$ and the conditional standard deviation, obtained as the square root of the conditional variance $\text{Var}\left(f_{e}^{\mathcal{R}_{e},\perp}\middle|\left\lVert{\bf p}_{\mathcal{R}_{ed}}^{\mathcal{R}_{e}}\right\rVert\right)$, in both scenarios.
+Visual inspection showcases a significant reduction in the expected normal interaction force in Scenario A, given the same virtual penetration commanded.
 Figure 8: Comparison between the conditional expectation of the interaction force in Scenario A and Scenario B.
-To evaluate the effectiveness of the force limitation strategy, we compared $E\left(A|B=b_{d}\right),\forall b_{d}\in\mathcal{B}_{d}$ and $\text{Var}(A|B=b_{d}),\forall b_{d}\in\mathcal{B}_{d}$ in both scenarios.
+To evaluate the effectiveness of the force limitation strategy, we compare $E\left(f_{e}^{\mathcal{R}_{e},\perp}\middle|\left\lVert{\bf p}_{\mathcal{R}_{ed}}^{\mathcal{R}_{e}}\right\rVert\right)$ and $\text{Var}\left(f_{e}^{\mathcal{R}_{e},\perp}\middle|\left\lVert{\bf p}_{\mathcal{R}_{ed}}^{\mathcal{R}_{e}}\right\rVert\right)$ in both scenarios.
 To test whether the conditional variances differ significantly, we perform Levene’s test and Bartlett’s test.
@@
 However, they do not provide any suggestion regarding the effectiveness of eye-hand coordination for the human operator.
-Therefore, in we provide a comprehensive qualitative assessment of the latter.
+Therefore, in the accompanying video [^1] This paper has supplementary downloadable material available at http://ieeexplore.ieee.org, provided by the authors.
+This includes a multimedia MOV format movie clip, showing the end-to-end system experimental demonstration.
+This material is 51.4 MB in size.
+we provide a comprehensive qualitative assessment of the latter.
 IV-D End-to-end system experimental demonstration
 We provide a demonstration of the end-to-end system’s operation in the accompanying video, where we showcase the execution of the dental scaling procedure on the three target teeth in Scenario A.
-In this section, we compare the outcomes of the execution of a dental scaling procedure in Scenario A, against those obtained in Scenario B.
-To avoid repetition, the analysis focuses on the outcomes of executing a dental scaling action on a single target tooth, the lower right first premolar.
-In Figure 10 the normal forces exerted by the FR in both scenarios are shown.
+In this section, we report the outcomes of the execution of a dental scaling procedure in Scenario A, highlighting two instants of the experiment (Figure 10), and compare them against those obtained in Scenario B (Figure 11).
+To avoid repetition, the analysis focuses on the outcomes of executing a dental scaling action on a single target tooth, the upper right second premolar.
+In Figure 10d and Figure 11b, the normal forces exerted by the FR in both scenarios are shown.
 A visual inspection showcases a significant reduction in interaction forces when the free-hand force limitation strategy is active, with a decrease of approximately $50\%$ in Scenario A, thereby confirming the validation presented in Section IV-B.
-The effectiveness of this strategy is made even more evident by Figure 11, which shows the norm of the references provided to the admittance controller.
+The effectiveness of this strategy is made even more evident by comparing Figure 10c and Figure 11a.
 Notably, in Scenario A, where the interaction forces are lower, the virtual penetration commanded to the controller is even greater than in Scenario B.
-Nevertheless, as shown in Figure 11a, the implementation of this strategy enables the scaling of the provided references, effectively reducing the contact forces and enhancing the system’s safety.
-(a) Contact forces in Scenario A.
-(b) Contact forces in Scenario B.
-Figure 10: Nomal forces in both scenarios.
-(a) Norm of the admittance controller references in Scenario A.
-(b) Norm of the admittance controller references in Scenario B.
-Figure 11: Norm of the admittance controller references in both scenarios.
-Finally, in Figure 12 the forces rendered on the HD in both scenarios are shown.
+Nevertheless, as shown in Figure 10c, the implementation of this strategy enables the scaling of the provided references, effectively reducing the contact forces and enhancing the system’s safety.
+[…]
+[…]
+(c) Norm of the admittance controller reference
+(d) Contact forces
+(e) Forces rendered on the HD
+Figure 10: Snapshots and outcomes of the dental scaling experiment in Scenario A.
+(a) Norm of the admittance controller reference
+(b) Contact forces
+(c) Forces rendered on the HD
+Figure 11: Outcomes of the dental scaling experiment in Scenario B.
+Finally, in Figure 10e and Figure 11c the forces rendered on the HD in both scenarios are shown.
 As expected, the results are consistent between the two scenarios, while both showing high similarity to exerted force profiles as well as slow dynamics.
 Such aspects enhance naturalness for the human operator.
-(a) Forces rendered on the HD in Scenario A.
-(b) Forces rendered on the HD in Scenario B.
-Figure 12: Forces rendered on the HD in both scenarios.
 V Discussions and conclusions
-Motivated by a lack of research on robotized free-hand dental procedures, this paper proposed a HBTS specifically designed to this aim.
+This paper proposed a FH-HBTS for free-hand dental procedures.
 The system features visual and haptic feedback, that are appropriately combined and coordinated to allow for a natural teleoperation experience.
 To improve the visibility of the intervention site, we designed a dedicated end-effector accommodating an eye-in-hand endoscopic camera, and ensuring compatibility with existing clinical tools.
-We improved the accuracy of the procedure by introducing human tremor filtering and dynamic scaling of velocities.
+We also improved the accuracy of the procedure by introducing human tremor filtering and dynamic scaling of velocities.
 By providing the eye-hand coordination and the TCP control, we enabled the human operator to perceive the motion correspondence between their movements and those of the end-effector, observed in the camera view, preserving manual dexterity and expertise.
@@
 The proposed strategy allows for accurate force limitation in any contact state and for any orientation of the interaction tool, without requiring any prior modeling of the intervention site.
-We believe that our HBTS is a necessary intermediate step for a broader adoption of robotics in dentistry, as it allows for accurate treatments, while preserving the operator’s cognitive and technical abilities.
+We believe that our FH-HBTS represents a crucial step towards a broader adoption of robotics in dentistry, as it allows for accurate treatments, while preserving the operator’s cognitive and technical abilities.
 With respect to manual procedures, our system allows to preserve operator’s ergonomics, track operator’s performance, and enable remote healthcare.
-We emphasize that the proposed system allows for the collection of haptic and motion data, which can be crucial during the training phase of new operators for the evaluation of learning metrics, i.e learning curves [55], or for the derivation of data-driven control policies [56] for autonomous robots.
-With our future work we intend to extend the system with autonomous patient motion compensation to further enhance the system’s safety.
-Likewise, we intend to conduct experiments with professionals and active dental tools both to confirm the system’s usability under more representative clinical conditions and to collect expert policies for (semi-)autonomous procedures.
+We emphasize that the proposed system allows to collect haptic and motion data.
+While this feature is crucial during the training phase of new operators for the evaluation of learning metrics, i.e learning curves [56], it is also central towards the deployment of data-driven control architectures that leverage demonstrations to synthesize control policies [57].
+Finally, these data can also be used to learn cost functions [58], and hence intentions, of experts fulfilling a given task.
+In the context of dental procedures, this means that data collected from experts executing the procedure can be used to train policies for autonomous robots.
+With our future work we plan to extend the FH-HBTS with autonomous patient motion compensation to further enhance the system’s safety, and to benchmark the performance against the da Vinci Surgical SystemTM.
+Also, we plan to conduct experiments with professionals and active dental tools both to confirm the system usability under more representative clinical conditions and to collect expert policies for (semi-)autonomous procedures.
 Acknowledgments

### v3

@@
 Robots can play a key role in improving procedural accuracy and safety, enhancing patient comfort, and reducing operator workload.
-However, robotic solutions for free-hand procedures remain limited or completely lacking, and their acceptance is still low.
+However, robotic solutions for free-hand procedures remain limited or completely lacking.
 To address this gap, we develop a haptic bilateral teleoperation system (HBTS) for free-hand dental procedures (FH-HBTS).
-The system includes a dedicated mechanical end-effector, compatible with standard clinical tools, and equipped with an endoscopic camera for improved visibility of the intervention site.
-By ensuring motion and force correspondence between the operator’s actions and the robot’s movements, monitored through visual feedback, we enhance the operator’s sensory awareness and motor accuracy.
-Furthermore, recognizing the need to ensure procedural safety, we limit interaction forces by scaling the motion references provided to the admittance controller based solely on measured contact forces.
+The system includes a mechanical end-effector, compatible with standard clinical tools, and equipped with an endoscopic camera for improved visibility of the intervention site.
+By ensuring motion and force correspondence between the operator’s and the robot’s actions, monitored through visual feedback, we enhance the operator’s sensory awareness and motor accuracy.
+Furthermore, to ensure procedural safety, we limit interaction forces by scaling the motion references provided to the admittance controller based solely on measured contact forces.
 This ensures effective force limitation in all contact states without requiring prior knowledge of the environment.
-The proposed FH-HBTS is validated in a dental scaling procedure using a dental phantom.
+The proposed FH-HBTS is validated both through a technical evaluation and an in-vitro pre-clinical study conducted on a dental model under clinically representative conditions.
 The results show that the system improves the naturalness, safety, and accuracy of teleoperation, highlighting its potential to enhance free-hand dental procedures.
-Index Terms:
-Haptics, teleoperation, interaction control, free-hand dental procedures, eye-hand coordination
 publicationid: pubid:
@@
 Traditionally, routine, prosthetic, periodontal, and some surgical dental procedures are performed free-hand.
-Routine procedures include the treatment of dental caries and endodontic therapy, characterized by the mechanical excavation of the damaged or infected dental structure, typically utilizing burs attached to a rotary dental handpiece, followed by the replacement of the lost structure with a synthetic filling material [2].
-Prosthetic procedures, on the other hand, focus on the replacement of missing or severely damaged teeth through the use of prostheses, such as inlays, crowns, or bridges.
-These procedures involve the preparation of the tooth, which entails the removal of hard tissue from the damaged or healthy tooth using rotary instruments mounted on dental handpieces [3].
-Periodontal procedures include scaling and root planing (SRP), which involve the removal of plaque and calculus from the surfaces of the teeth and the smoothing of root surfaces, typically employing manual hand instruments and inserts mounted on sonic or ultrasonic handpieces [4].
-Finally, although robotic computer-assisted implant surgery (r-CAIS) have shown a significant increase in the accuracy of dental implant placement [5], many operators still prefer the free-hand method for dental implantology [6].
+Routine procedures include the treatment of dental caries and endodontic therapy, while prosthetic procedures focus on the replacement of teeth through the use of prostheses.
+These procedures involve the preparation of the tooth, which entails the removal of hard tissue using rotary instruments mounted on dental handpieces [43].
+Periodontal procedures include scaling and root planing (SRP), i.e., the removal of plaque and calculus from the teeth and the smoothing of root surfaces, typically employing hand instruments and inserts mounted on sonic or ultrasonic handpieces [53].
+Finally, although robotic computer-assisted implant surgery (r-CAIS) have shown a significant increase in the accuracy of dental implant placement, many operators still prefer the free-hand method for dental implantology [2].
+Each of these procedures requires accuracy, extensively trained manual skills, and dexterity.
 Figure 1: Manual (left) vs teleoperated (right) dental procedures (pictorial views).
-Each of these procedures, even the simplest ones, require accuracy, extensively trained manual skills and dexterity.
 Remark 1.
@@
 The main causes of these disorders can be traced back to an incorrect application of ergonomics, such as awkward postures, repetitive and prolonged movements, use of small or narrow tool handles, exertion of forces on hard or sharp objects, and exposure to vibrations [24].
-Hence, when training new operators, much emphasis is placed on learning skills related to space utilization and positioning of surgical and non-surgical instruments, e.g., mirrors, to overcome occlusions, while maintaining healthy and comfortable posture [12].
-In addition to this, to ensure the safety and effectiveness of the procedure, new operators must acquire skills related to the correct amount of pressure to apply to the site of interest, depending on the type of tool used.
+Hence, when training new operators, much emphasis is placed on learning skills related to space utilization and positioning of surgical and non-surgical instruments, e.g., mirrors, to overcome occlusions, while maintaining healthy and comfortable postures [15].
+In addition, to ensure the safety and effectiveness of the procedure, new operators must acquire skills related to the correct amount of pressure to apply to the site of interest, depending on the type of tool used.
 Remark 2.
 The sense of touch plays a key role in achieving successful and efficient procedures.
-This is also apparent from emerging trends in teaching, which make use of virtual environments and haptic devices (HD)s for training new operators [13, 14].
-Robots can achieve accuracy, precision, efficiency beyond human capabilities.
-For this reason, the role of robots is to improve dental procedures, both in terms of patient satisfaction, i.e.
-broader access to care, reduced treatment time, fast recovery, and in terms of operator comfort, i.e., reduced workload intensity, increased visibility of the surgical field, improved ergonomics [15], as shown in Figure 1.
+This is also apparent from emerging trends in teaching, which make use of virtual environments and haptic devices (HDs) for training new operators [37, 59].
+In this context, the role of robots is to improve dental procedures, in terms of both patient satisfaction, i.e., broader access to care, reduced treatment time, and fast recovery, and operator comfort, i.e., reduced workload intensity, increased visibility of the surgical field, and improved ergonomics [1], as shown in Figure 1.
 I-A Related works
-Haptic teleoperations are a technology that enables the remote execution of complex tasks, employing a robotic system.
+Haptic teleoperations are an enabling technology for the remote execution of complex tasks, employing a robotic system.
 Remark 3.
@@
 By providing multiple types of sensory information, i.e.
-visual and haptic feedback, HBTSs enable the human operator to experience telepresence, i.e.
-the perception of interacting with the remote environment as if it were directly encountered [16].
+visual and haptic feedback, HBTSs enable the human operator to experience telepresence [45].
 Furthermore, by integrating different haptic feedback and guidance modalities [67, 30], and shared control approaches [12], they contribute to reducing errors and enhance the efficiency, accuracy and safety of procedures, while also improving operator comfort and confidence.
-For these reasons, such technology is applied in multiple domains, such as telesurgery, space exploration, and the handling of hazardous materials in nuclear or underwater environments.
-However, in the context of dentistry, the adoption of such systems is still very limited.
-The majority of research in robotics for dentistry is focused on oral implantology [20].
-Robotic oral implant surgery offers significant benefits in terms of implant accuracy in partially [21] or totally edentulous patients, also compared to computer-assisted implant surgery (s-CAIS) systems [22].
-Better placement accuracy, however, can be achieved at the expense of surgical efficiency, as shown in [23].
-Such technologies also enable flapless dental implant placement, speeding up recovery time, reducing postoperative pain, and enhancing patient comfort [24].
-[25] shows how a telerobotic-assisted drilling system can enhance the accuracy of the human operator while preserving awareness of the procedure through an effective transmission of cutting forces and vibrations.
-On the other hand, the state-of-the-art shows considerably less research effort in other free-hand procedures [26].
-In the context of prosthetic procedures, robotics can enhance the precision of manual grinding using a turbine drill during tooth preparation [27] and can also compensate for hand tremors and the lack of tactile feedback in laser ablation [28, 29], improving accuracy [30] and efficiency [31].
-In endodontics, robotics can automate and enhance precision and efficiency in root canal treatments [32, 33], outperforming current approaches also in effectively addressing biological and mechanical traits associated with biofilm recalcitrance and reinfection [34].
-In periodontal procedures, robotic systems have the potential to increase accuracy, reducing the risk of errors and trauma to tissues by optimizing applied forces, as with other free-hand procedures.
-They can also improve the ergonomics for the operator, and comfort for the patient by speeding up recovery time.
+For these reasons, such technology is consolidated in, e.g., telesurgery, space exploration, and the handling of hazardous materials in nuclear or underwater environments.
+However, in the context of dentistry, its adoption is still very limited.
+The majority of research in robotics for dentistry is focused on oral implantology [6], mainly because robotic implant surgery offers significant benefits in terms of implant accuracy in partially [65] or totally edentulous patients.
+Such technology also enables flapless implant placement, accelerating recovery time, reducing postoperative pain, and enhancing patient comfort [58].
+[28] shows that a telerobotic-assisted drill can enhance accuracy while preserving awareness through an effective transmission of cutting forces and vibrations.
+On the other hand, the state-of-the-art shows considerably less research effort in other free-hand procedures, a trend also highlighted in existing reviews of dental robotics [3], which report a strong imbalance in favor of implantology.
+Despite this imbalance, the literature reports evidence of the potential benefits of robotic technologies in free-hand dental procedures.
+Such evidence is analyzed in the following and summarized in the supplementary materials [^1] This paper has supplementary downloadable material available at http://ieeexplore.ieee.org, provided by the authors.
+This includes a supplementary PDF format information file and a multimedia MOV format movie clip, showing the end-to-end system experimental demonstration.
+This material is 78.5 MB in size..
+The analysis focuses on robotic systems explicitly developed for dental procedures and directly interacting with the oral environment, aiming at supporting or automating clinical free-hand procedures, while excluding works specifically addressing oral implantology.
+Only approaches supported by experimental, in vitro, or preliminary validation are considered, whereas purely virtual studies or robotic solutions not specifically tailored to dental procedures are not included.
+In the context of prosthetic procedures, robotics can enhance the precision of manual grinding using a turbine drill during tooth preparation [27] and can compensate for hand tremors and the lack of tactile feedback in laser ablation [62, 63].
+This can improve accuracy [35] and efficiency [68], and can enable automated tooth arrangement [70] and preparation for veneers and full crowns [40, 69].
+In endodontics, robotics can automate and enhance precision and efficiency in root canal treatments [13, 14], outperforming current approaches also in effectively addressing biological and mechanical traits associated with biofilm recalcitrance and reinfection [26], and can support visual-guided tool navigation [23].
+In periodontal procedures, robots have the potential to increase accuracy, reducing the risk of errors and trauma to tissues by optimizing applied forces, as with other free-hand procedures.
+They can also improve the ergonomics for the operator, and comfort for the patient by reducing recovery time.
 However, for such procedures, robotic technologies are still lacking [33], and there is no empirical evidence of such advantages.
+Existing systems addressing force-regulated interaction with teeth are limited to adjacent tasks, such as tooth-brushing assistance [4], and are not designed for direct periodontal procedures.
 Thus, although robotics shows promise in dental practice, several limitations remain.
 First, the quality of research is still insufficient in terms of clinical validation and technology readiness level (TRL) [61].
-This is heavily influenced by the high costs of these robotic systems and the need to ensure their compatibility with existing clinical tools [37].
-Second, the most widely used robots are fully autonomous, which require the development of state-of-the-art algorithms and advanced capabilities in perception, recognition, and exception handling [37].
+This is heavily influenced by the high costs of these systems and the need to ensure their compatibility with existing clinical tools [66].
+Second, the most widespread robots are autonomous, which require the development of advanced capabilities in perception, recognition, and exception handling [66].
 In this context, the presence of a human is still necessary not only as a supervisor but also as an operator.
 In light of Remark 2 and Remark 3, and considering the impact on the operator acceptance [32], haptic teleoperations represent a viable solution.
-Moreover, the HBTSs would allow the acquisition of human control strategies directly exercised in the robot state space, simplifying the process of transferring the knowledge from the human to the robot [39] and enabling the effective and objective evaluation of operator skills during training [40].
+Moreover, HBTSs would allow the acquisition of human control strategies directly exercised in the robot state space, simplifying the process of transferring the knowledge from the human to the robot [46] and enabling the effective and objective evaluation of operator skills during training [31].
+However, as emerges from the literature analysis above, no robotic solutions based on haptic bilateral teleoperation have been reported for free-hand tasks.
 Finally, research tends to prioritize oral implantology over other dental procedures, despite the high number of daily free-hand routine procedures and the less invasiveness, entailing better patient acceptance [38].
-Considering these limitations, [26] recommends to shift the focus from oral implantology procedures towards less invasive ones, which are better suited as forerunners in the field, and also to invest in educational robotics research in academia as a valuable tool to overcome the barrier of acceptance of robotic systems among future dentists.
+Considering these limitations, [3] recommends to shift the focus from oral implantology towards less invasive procedures, which are better suited as forerunners in the field, and also to invest in educational robotics research as a valuable tool to overcome the barrier of acceptance of robotic systems among future dentists.
 I-B Contributions
-Motivated by the limitations or complete absence of robotic solutions for the most repetitive free-hand dental procedures, and in light of Remark 3, in this paper we present a HBTS for free-hand dental procedures (FH-HBTS), including:
-  • a dedicated mechanical end-effector, compatible with existing clinical tools and featuring an endoscopic camera, to improve visibility at the intervention site;
-  • a virtualized force feedback, consistent with the interaction observed through the visual feedback of the eye-in-hand camera, which enhances the haptic experience for the operator;
-  • a control system that ensures procedural safety, improves accuracy, and preserves manual dexterity, while providing motion correspondence between the operator’s movements and those of the end-effector, as observed in the camera view.
-To the best of the authors’ knowledge, this is the only HBTS designed for free-hand dental procedures. Key contributions are:
-  • a control system to ensure accurate control of the tool center point (TCP) and to guarantee motion correspondence, using a novel eye-hand coordination strategy;
-  • a force limitation strategy that makes no previous assumption or estimation of the environment geometry, therefore accounting for the challenges outlined in Remark 1;
-  • eye-hand force correspondence provided with saturation and filtering via a deadband function;
-  • TRL-4 validation of the FH-HBTS on a free-hand procedure, consisting of dental scaling.
+Motivated by the limitations or complete absence of robotic solutions for the most repetitive free-hand dental procedures, and in light of the advantages introduced by haptic teleoperations, as in Remark 3, we propose a HBTS for free-hand dental procedures (FH-HBTS).
+The proposed system features a custom designed end-effector, compatible with existing clinical tools, and a dedicated haptic control system that ensures procedural safety, improves accuracy, and preserves manual dexterity.
+Therein, key contributions are: (1) an eye-hand coordination controller providing both motion and force correspondence between the HD’s workspace and the video stream coming from the eye-in-hand camera; with respect to existing HBTSs, such mapping is bilateral and drift-free, even after prolonged teleoperation sessions; (2) a force limitation strategy effectively reducing interaction forces; with respect to existing HBTSs, this strategy needs no previous assumption or online estimation of the environment geometry, therefore accounting for the challenges outlined in Remark 1; notably, this strategy is part of the onboard robot controller and is therefore compatible with passivity-based teleoperations; (3) an integrated TRL-5 system for free-hand dental procedures, validated on SRP, demonstrating technical feasibility and clinical suitability.
+In light of the inclusion and exclusion criteria defined in Section I-A, and also reported in the supplementary materials, to the best of the authors’ knowledge, this is the first HBTS specifically complying with the special requirements of safety, accuracy, and dexterity characterizing free-hand dental procedures.
 II Preliminaries
@@
 ${\bf T}_{b}^{a}=\begin{pmatrix}{\bf R}_{b}^{a}&{\bf p}_{b}^{a}\\ {\bf 0}^{\top}&1\end{pmatrix}$, (1)
-where ${\bf R}_{b}^{a}\in SO(3)$ is the rotation matrix from frame $a$ to frame $b$, and ${\bf p}_{b}^{a}\in\mathbb{R}^{3\times 1}$ represents $b$’s origin in frame $a$.
-We let ${\bf x}_{b}^{a}$ be the pose of the frame $b$ with respect to the frame $a$, corresponding to ${\bf T}_{b}^{a}$, in the form:
-${\bf x}_{b}^{a}=\begin{bmatrix}{\bf p}_{b}^{a}\\ {\bf q}_{b}^{a}\end{bmatrix},$ (2)
-where ${\bf q}_{b}^{a}\in\mathbb{H}_{1}$ is the quaternion associated with ${\bf R}_{b}^{a}$.
+where ${\bf R}_{b}^{a}\in SO(3)$ is the rotation matrix from frame $a$ to frame $b$, with ${\bf R}_{a}^{b}=({\bf R}_{b}^{a})^{-1}\equiv({\bf R}^{a}_{b})^{\top}$, and ${\bf p}_{b}^{a}\in\mathbb{R}^{3\times 1}$ represents $b$’s origin in frame $a$.
+We let ${\bf x}_{b}^{a}=\begin{bmatrix}({\bf p}_{b}^{a})^{\top},({\bf q}_{b}^{a})^{\top}\end{bmatrix}^{\top}$ be the pose of the frame $b$ with respect to the frame $a$, corresponding to ${\bf T}_{b}^{a}$, where ${\bf q}_{b}^{a}\in\mathbb{H}_{1}$ is the quaternion associated with ${\bf R}_{b}^{a}$.
 We denote by ${\bf R}_{\bf a}^{\mathcal{F}}(\alpha)\in SO(3)$ the elementary rotation of an angle $\alpha$ about a generic axis $\bf a$ of a frame $\mathcal{F}$.
-We denote by ${\bf v}^{a}_{b}\in\mathbb{R}^{6\times 1}$ the twist of the frame $b$ with respect to the frame $a$, in the form:
-${\bf v}^{a}_{b}=\begin{bmatrix}\dot{\bf p}^{a}_{b}\\ {\boldsymbol{\omega}}^{a}_{b}\end{bmatrix},$ (3)
-where $\dot{\bf p}^{a}_{b},{\boldsymbol{\omega}}^{a}_{b}$ are the linear and angular velocities, respectively.
+We denote by ${\bf v}^{a}_{b}=\begin{bmatrix}(\dot{\bf p}^{a}_{b})^{\top},({\boldsymbol{\omega}}^{a}_{b})^{\top}\end{bmatrix}^{\top}\in\mathbb{R}^{6\times 1}$ the twist of the frame $b$ with respect to the frame $a$, where $\dot{\bf p}^{a}_{b},{\boldsymbol{\omega}}^{a}_{b}\in\mathbb{R}^{3\times 1}$ are the linear and angular velocities, respectively.
+We let ${\bf h}^{a}=\begin{bmatrix}({\bf f}^{a})^{\top},({\bf m}^{a})^{\top}\end{bmatrix}^{\top}\in\mathbb{R}^{6\times 1}$ be the wrench with respect to the frame $a$, where ${\bf f}^{a},{\bf m}^{a}\in\mathbb{R}^{3\times 1}$ are the forces and momenta, respectively.
+${\bf I}_{n}$ indicates the $n\times n$ identity matrix.
 II-B Haptic device and robot reference frames
-The frames used by the HD and the robot are reported in Figure 2. With regard to the HD, two frames are introduced:
-  • $\mathcal{H}_{b}$ is the base reference frame, in which the HD’s stylus poses and the feedback forces are expressed;
-  • $\mathcal{H}_{e}$ is the HD’s stylus frame.
-Regarding the robot, four frames are defined:
-  • $\mathcal{R}_{b}$ is the base reference frame, in which motion references are expressed;
-  • $\mathcal{R}_{e}$ is the robot’s end-effector frame (computed from forward kinematics of joint measurements), corresponding to the TCP;
-  • $\mathcal{R}_{ed}$ is the desired robot’s end-effector frame, input to admittance control;
-  • $\mathcal{R}_{c}$ is the compliant frame, output by the admittance control law.
+The frames used by the HD and the robot are reported in Figure 2.
+With regard to the HD, two frames are introduced: $\mathcal{H}_{b}$ is the base reference frame, in which the HD’s stylus poses and the feedback forces are expressed; $\mathcal{H}_{e}$ is the HD’s stylus frame.
+Regarding the robot, four frames are defined: $\mathcal{R}_{b}$ is the base reference frame, in which motion references are expressed; $\mathcal{R}_{e}$ is the robot’s end-effector frame (computed from forward kinematics of joint measurements), corresponding to the tool center point (TCP); $\mathcal{R}_{ed}$ is the desired robot’s end-effector frame, input to admittance control; $\mathcal{R}_{c}$ is the compliant frame, output by the admittance control law.
 We note that, after admittance control, motion control guarantees that $\mathcal{R}_{e}$ tracks $\mathcal{R}_{c}$, as discussed next.
@@
 III Methodology
-The FH-HBTS we propose, shown in Figure 3, is based on a position-force architecture [42], coupling the leader (LR) and follower (FR) robots.
-The LR is an impedance HD, sensing the pose of the human operator and rendering haptic feedback according to the remote interaction between the FR and the environment.
+The FH-HBTS we propose, shown in Figure 3, is based on a position-position architecture [39], coupling the leader (LR) and follower (FR) robots by means of a spring–damper, which is a passive element.
+Therefore, in the absence of communication delays, this architecture has the advantage of being intrinsically stable [5, 25].
+A formal proof of absolute stability can be found in [47].
+In the proposed design, we assume negligible time delays and packet losses and, accordingly, do not integrate passivity-based control approaches [42].
+This choice is motivated both by the validity of this assumption in healthcare facilities and by the conservative nature of passivity-based methods, which are known to worsen transparency and system usability in order to ensure stability under adverse network conditions [52].
+Indeed, in precision free-hand dental procedures, this trade-off is generally undesirable, as it may impair fine force discrimination and the naturalness of the interaction.
+Nevertheless, the management of communication delays and packet losses represents an orthogonal direction to the one investigated in this work.
+Indeed, under non-ideal network conditions, the proposed architecture can be readily extended to enforce time-domain passivity of the communication channel, for instance by adopting a TDPA-based approach [7], as demonstrated in Section IV-B.
+The LR is an impedance HD, sensing the pose of the operator and rendering haptic feedback according to the remote interaction between the FR and the environment.
 The HD is controlled by a leader controller, featuring an eye-hand coordination controller (EHCC) and a haptic feedback controller (HFC).
-The former is responsible for capturing the HD’s stylus poses, filtering the human tremors, and generating the corresponding robot end-effector poses, so that the robot follows the stylus motion in coordination with the visual feedback.
+The former is responsible for capturing the HD’s stylus poses, filtering human tremors, and generating the corresponding robot end-effector poses, so that the robot follows the stylus motion in coordination with the visual feedback.
 The latter is responsible for rendering the force feedback on the HD to let the human operator perceive the remote environment.
 Both controllers record data throughout the entire procedure.
-The FR is a position-controlled robot, instructed to track the HD through an admittance controller, which provides the latter with a configurable degree of compliance on its target pose with respect to the environment.
-Such controller has three key features.
-First, it ensures the stability of the FR, even in case of loss of communication, since, in case of a lack of reference from the EHCC, the controller keeps following the last one received.
-Second, it avoids damage to the arm structure and manipulated environment, since it adapts the received references to ensure compliant interaction with the surroundings.
-Finally, it limits the interaction forces by scaling the received motion references through a free-hand force limitation strategy embedded in the controller itself.
+The FR is a position-controlled robot, instructed to track the HD through an admittance controller, which provides a configurable degree of compliance on its target pose with respect to the environment and improves safety through dedicated mechanisms and emergency protocols.
 With reference to Figure 3, depicting the complete system, the following sections examine its individual subcomponents in detail.
-In Section III-A and Section III-C, a detailed description of the objectives and operation of the EHCC and HFC controllers is provided.
-In Section III-B the robot interaction control is described.
+Section III-A and Section III-C provide a detailed description of the objectives and operation of the EHCC and HFC controllers.
+Section III-B describes the robot interaction control.
+Finally, Section III-D presents the design of the mechanical end-effector for the FR.
 Figure 3: Haptic bilateral teleoperation system scheme.
@@
 This approach enables the operator to maintain a consistently centered view of the operative field, offering a significant advantage over eye-to-hand camera configurations, where the view may be off-center with respect to the area of interest or obstructed by physical obstacles.
-In this context, the human operator is naturally guided by the video stream transmitted by the camera to provide commands to the FR via the HD.
-However, in the eye-in-hand configuration, a rotation of the tool (unavoidable, to preserve dexterity) causes a rotation of the viewing angle.
-This creates a mismatch between the movements expected by the operator and those actually executed by the robot in the camera view, compromising the accuracy of the task.
-The da Vinci Surgical SystemTM, which requires the insertion of multiple tools to implement an eye-to-hand strategy, thereby increasing the bulk of the equipment and the overall costs, provides the motion correspondence between the operator’s movements and those of the end-effector, observed in the camera view, by projecting the endoscope image above the surgeon’s hands by means of mirrored overlay optics [44].
-In our eye-hand coordination controller, we provide the motion correspondence at the command level, by creating a coupling between the HD’s stylus and the robot’s TCP, and appropriately converting the motion of the former into the motion of the FR.
+In this context, the human operator relies on the video of the camera to command the FR via the HD.
+However, in the eye-in-hand configuration, a rotation of the tool (unavoidable, to preserve dexterity) causes a rotation of the viewing angle, whose lack of compensation results in a loss of both motion and force correspondence between the movements and perceived interaction expected by the operator and those actually executed by the robot in the camera view.
+This highlights the importance of an appropriate coupling.
+The literature shows that this coupling can be achieved through simple kinematic transformations and scaling strategies [44].
+Common mapping strategies include position control, workspace drift control [11], and rate control [51] which define how the operator’s motions are translated into robot commands, complemented by force scaling [44] to map the interaction forces into the range of forces that the HD can exert.
+However, these methods generally neglect viewing angle effects, relying on display-level compensation and thus compromising force correspondence.
+Indeed, the da Vinci Surgical SystemTM, which requires the insertion of multiple tools to implement an eye-to-hand strategy, thereby increasing bulk and cost of the equipment, provides the motion correspondence by projecting the endoscope image above the surgeon’s hands by means of mirrored overlay optics [18].
+In a similar manner, in eye-in-hand cases, motion correspondence is typically achieved by aligning the visual feedback with the operator’s control frame through online compensation of camera–control frame misalignment in the displayed view [64].
+To overcome this limitation, we introduce an eye–hand coordination mapping preserving both motion and force correspondence.
+Within the proposed EHCC, this mapping is implemented through a view-consistent transfer of the HD twist to the FR and incorporates dynamic scaling factors to improve control accuracy according to task requirements.
 To enhance the naturalness of the teleoperation, a rotational coupling is also established between the HD’s stylus and the FR’s TCP.
-This includes defining the desired relative rotation ${\bf R}_{\mathcal{H}_{e}}^{\mathcal{R}_{e}}$, and performing manually the alignment.
+This includes defining the desired relative rotation ${\bf R}_{\mathcal{R}_{e}}^{\mathcal{H}_{e}}$, and aligning manually.
 This is done by the human operator during what we indicate as the engagement phase.
 The latter is presented in Section III-A1, while the human tremors filtering and the eye-hand coordination mapping, implementing the motion correspondence, are discussed in Section III-A2 and Section III-A3, respectively.
+The eye-hand coordination implementing the force correspondence is described in Section III-C.
 III-A1 Engagement phase
 During the engagement phase, the EHCC retrieves the homogeneous transformation matrix of the HD’s stylus ${\bf T}_{\mathcal{H}_{e}}^{\mathcal{H}_{b}}(k)$, and converts it into a robot’s end-effector homogeneous transformation matrix with the following kinematic equations:
-${\bf R}_{\mathcal{R}_{ed}}^{\mathcal{R}_{b}}(k)$ $={\bf R}_{\mathcal{H}_{b}}^{\mathcal{R}_{b}}{\bf R}_{\mathcal{H}_{e}}^{\mathcal{H}_{b}}(k)\left({\bf R}_{\mathcal{H}_{e}}^{\mathcal{R}_{e}}\right)^{\top},$ (4)
+${\bf R}_{\mathcal{R}_{ed}}^{\mathcal{R}_{b}}(k)$ $={\bf R}_{\mathcal{H}_{b}}^{\mathcal{R}_{b}}{\bf R}_{\mathcal{H}_{e}}^{\mathcal{H}_{b}}(k){\bf R}_{\mathcal{R}_{e}}^{\mathcal{H}_{e}},$ (2)
 ${\bf p}_{\mathcal{R}_{ed}}^{\mathcal{R}_{b}}(k)$ $={\bf p}_{\mathcal{R}_{e}}^{\mathcal{R}_{b}}(k),$ (3)
@@
 The obtained frame $\mathcal{R}_{ed}$ is real-time displayed in a visualization tool, together with the actual robot end-effector one $\mathcal{R}_{e}$, as shown in Figure 4, so that the human operator, by acting on $\mathcal{H}_{e}$, can align the frame $\mathcal{R}_{ed}$ with the frame $\mathcal{R}_{e}$.
-Once the alignment is accomplished, i.e.
-$\left[{\bf R}_{\mathcal{R}_{e}}^{\mathcal{R}_{b}}(k)\right]^{\top}{\bf R}_{\mathcal{R}_{ed}}^{\mathcal{R}_{b}}(k)={\bf I}_{3}$, where ${\bf I}_{3}\in SO(3)$ is the identity matrix, the engagement phase ends, i.e.
-HD and FR are aligned, and the human operator can begin teleoperation.
+Once the alignment is accomplished, i.e., ${\bf R}_{\mathcal{R}_{b}}^{\mathcal{R}_{e}}(k){\bf R}_{\mathcal{R}_{ed}}^{\mathcal{R}_{b}}(k)={\bf I}_{3}$, the engagement phase ends, i.e., HD and FR are aligned, and the human operator can begin teleoperation.
 Figure 4: Engagement phase supported by the visualization tool.
@@
 At each control cycle, the EHCC retrieves the actual pose ${\bf x}_{\mathcal{H}_{e}}^{\mathcal{H}_{b}}(k)$ of the HD’s stylus, corresponding to ${\bf T}_{\mathcal{H}_{e}}^{\mathcal{H}_{b}}(k)$.
-Exploiting such a pose to generate a command for the robot would result in transferring the human hand motion tremors, further accentuated by the control frequency.
+Exploiting such a pose to generate a command for the robot would result in transferring the human hand tremors, further accentuated by the control frequency.
 Therefore, ${\bf x}_{\mathcal{H}_{e}}^{\mathcal{H}_{b}}(k)$ is filtered online by a moving average filter for poses.
-At the $k$-th control cycle, the filtered position $\overline{{\bf p}}_{\mathcal{H}_{e}}^{\mathcal{H}_{b}}(k)$ is obtained from the cumulative sum ${\bf s}(k)$ of samples within a window of fixed size $n$:
-$\overline{{\bf p}}_{\mathcal{H}_{e}}^{\mathcal{H}_{b}}(k)=\frac{{\bf s}(k)}{m}$ (6)
-where $m\leq n$ is the number of samples in the moving average window at the $k$-th control cycle, and the cumulative sum ${\bf s}(k)$ is:
-${\bf s}(k)=\begin{cases}{\bf s}(k-1)+{\bf p}_{\mathcal{H}_{e}}^{\mathcal{H}_{b}}(k),&\text{if}~{}k<n;\\ {\bf s}(k-1)+{\bf p}_{\mathcal{H}_{e}}^{\mathcal{H}_{b}}(k)-{\bf p}_{\mathcal{H}_{e}}^{\mathcal{H}_{b}}(k-n),&\text{otherwise.}\end{cases}$ (7)
-The filtered quaternion $\overline{\bf q}_{\mathcal{H}_{e}}^{\mathcal{H}_{b}}(k)$ is the eigenvector associated with the maximum eigenvalue of the ${\bf M}(k)$ matrix, of proper size [45], defined as:
-${\bf M}(k)=\frac{{\bf c}(k)}{m},$ (8)
-where ${\bf c}(k)$ is the cumulative sum of samples within a window of fixed size $n$, updated at each control cycle:
-${\bf c}(k)=\begin{cases}{\bf c}(k-1)+{\bf q}_{\mathcal{H}_{e}}^{\mathcal{H}_{b}}(k)\left[{\bf q}_{\mathcal{H}_{e}}^{\mathcal{H}_{b}}(k)\right]^{\top},&\text{if}~{}k<n;\\ {\bf c}(k-1)+{\bf q}_{\mathcal{H}_{e}}^{\mathcal{H}_{b}}(k)\left[{\bf q}_{\mathcal{H}_{e}}^{\mathcal{H}_{b}}(k)\right]^{\top}+&\\ \quad-{\bf q}_{\mathcal{H}_{e}}^{\mathcal{H}_{b}}(k-n)\left[{\bf q}_{\mathcal{H}_{e}}^{\mathcal{H}_{b}}(k-n)\right]^{\top},&\text{otherwise.}\end{cases}$ (9)
+At the $k$-th control cycle, the filtered position is obtained as $\overline{{\bf p}}_{\mathcal{H}_{e}}^{\mathcal{H}_{b}}(k)={\bf s}(k)/m$, where $m\leq n$ is the number of collected samples at the $k$-th control cycle in the moving average window of size $n$, and ${\bf s}(k)$ is their cumulative sum:
+${\bf s}(k)={\bf s}(k-1)+{\bf p}_{\mathcal{H}_{e}}^{\mathcal{H}_{b}}(k)-\begin{cases}0,&\text{if}~k<n;\\ {\bf p}_{\mathcal{H}_{e}}^{\mathcal{H}_{b}}(k-n),&\text{otherwise.}\end{cases}$ (4)
+The filtered quaternion $\overline{\bf q}_{\mathcal{H}_{e}}^{\mathcal{H}_{b}}(k)$ is the eigenvector associated with the maximum eigenvalue of ${\bf M}(k)={\bf C}(k)/m$ [36], where ${\bf C}(k)$ is the cumulative sum of samples within the window, updated at each control cycle:
+${\bf C}(k)=\begin{cases}{\bf C}(k-1)+{\bf q}_{\mathcal{H}_{e}}^{\mathcal{H}_{b}}(k)\bigl[{\bf q}_{\mathcal{H}_{e}}^{\mathcal{H}_{b}}(k)\bigr]^{\top},&\text{if}~k<n;\\ {\bf C}(k-1)+{\bf q}_{\mathcal{H}_{e}}^{\mathcal{H}_{b}}(k)\bigl[{\bf q}_{\mathcal{H}_{e}}^{\mathcal{H}_{b}}(k)\bigr]^{\top}+&\\ \quad-{\bf q}_{\mathcal{H}_{e}}^{\mathcal{H}_{b}}(k-n)\bigl[{\bf q}_{\mathcal{H}_{e}}^{\mathcal{H}_{b}}(k-n)\bigr]^{\top},&\text{otherwise.}\end{cases}$ (5)
 Keeping the cumulative sums ${\bf s}(k)$ and ${\bf C}(k)$ enables achieving $\mathcal{O}(1)$ computational complexity.
 III-A3 Eye-hand coordination
-In a HBTS, coupling two robots with different kinematics requires the use of mapping and scaling techniques.
-The literature shows that this coupling can be achieved by converting the motion of the HD into that of the FR through simple kinematic transformations, and appropriately scaling it to either expand the workspace size of the HD or enhance control accuracy [46].
-In the eye-in-hand configuration, converting quantities between the two spaces without considering the viewing angle would lead to a loss of motion correspondence between the movements expected by the operator and those actually executed by the robot in the camera view.
-Therefore, in our EHCC, we transfer the HD’s twist to the FR through an eye-hand coordination mapping and introduce dynamic scaling factors to improve control accuracy based on the task requirements.
 Let $\Delta\overline{{\bf p}}_{\mathcal{H}_{e}}^{\mathcal{H}_{b}}(k)$ be the HD’s displacement:
 $\Delta\overline{{\bf p}}_{\mathcal{H}_{e}}^{\mathcal{H}_{b}}(k)=\overline{{\bf p}}_{\mathcal{H}_{e}}^{\mathcal{H}_{b}}(k)-\overline{{\bf p}}_{\mathcal{H}_{e}}^{\mathcal{H}_{b}}(k-1),$ (6)
-and $\overline{{\bf R}}^{\mathcal{H}_{e}(k-1)}_{\mathcal{H}_{e}(k)}$ the relative rotation:
-$\overline{{\bf R}}^{\mathcal{H}_{e}(k-1)}_{\mathcal{H}_{e}(k)}=\begin{bmatrix}r_{11}&r_{12}&r_{13}\\ r_{21}&r_{22}&r_{23}\\ r_{31}&r_{32}&r_{33}\end{bmatrix},$ (11)
-obtained from the quaternion $\overline{\bf q}_{\mathcal{H}_{e}(k)}^{\mathcal{H}_{e}(k-1)}$:
-$\overline{\bf q}_{\mathcal{H}_{e}(k)}^{\mathcal{H}_{e}(k-1)}=\left[\overline{\bf q}_{\mathcal{H}_{e}}^{\mathcal{H}_{b}}(k-1)\right]^{\top}\overline{\bf q}_{\mathcal{H}_{e}}^{\mathcal{H}_{b}}(k).$ (12)
+and $\overline{{\bf R}}^{\mathcal{H}_{e}(k-1)}_{\mathcal{H}_{e}(k)}$ the relative rotation obtained from
+$\overline{\bf q}_{\mathcal{H}_{e}(k)}^{\mathcal{H}_{e}(k-1)}=\bigl[\overline{\bf q}_{\mathcal{H}_{e}}^{\mathcal{H}_{b}}(k-1)\bigr]^{\top}\overline{\bf q}_{\mathcal{H}_{e}}^{\mathcal{H}_{b}}(k).$ (7)
 At each control cycle, the EHCC computes the HD’s stylus twist:
-${\bf v}_{\mathcal{H}_{e}}^{\mathcal{H}_{b}}(k)=\begin{bmatrix}\dfrac{\Delta\overline{{\bf p}}_{\mathcal{H}_{e}}^{\mathcal{H}_{b}}(k)}{T_{s}}\\[8.5359pt] \dfrac{\vartheta{\bf r}}{T_{s}}\end{bmatrix},$ (13)
-where $T_{s}$ is the control period and ${\bf r}$ and $\vartheta$ are the axis and the angle corresponding to $\overline{{\bf R}}^{\mathcal{H}_{e}(k-1)}_{\mathcal{H}_{e}(k)}$, respectively:
-$\vartheta$ $=\cos^{-1}{\left(\dfrac{r_{11}+r_{22}+r_{33}-1}{2}\right)},$ (14)
-${\bf r}$ $=\frac{1}{2\sin{\vartheta}}\begin{bmatrix}r_{32}-r_{23}\\ r_{13}-r_{31}\\ r_{21}-r_{12}\end{bmatrix}.$ (15)
+${\bf v}_{\mathcal{H}_{e}}^{\mathcal{H}_{b}}(k)=T_{s}^{-1}\begin{bmatrix}\Delta\overline{{\bf p}}_{\mathcal{H}_{e}}^{\mathcal{H}_{b}}(k)^{\top},&\vartheta{\bf r}^{\top}\end{bmatrix}^{\top}$, (8)
+where $T_{s}\in\mathbb{R}_{+}$ is the control period, and ${\bf r}\in\mathbb{R}^{3\times 1}$ and $\vartheta\in\mathbb{R}$ are the axis-angle representation corresponding to $\overline{{\bf R}}^{\mathcal{H}_{e}(k-1)}_{\mathcal{H}_{e}(k)}$, respectively.
 From (8), the corresponding desired FR’s end-effector twist is obtained as:
-$\begin{split}{\bf v}_{\mathcal{R}_{ed}}^{\mathcal{R}_{b}}(k)=&\left[{\bf I}_{2}\otimes{\bf R}_{\mathcal{R}_{e}}^{\mathcal{R}_{b}}(k)\right]{\boldsymbol{\Lambda}}\left[{\bf I}_{2}\otimes{\bf R}_{\mathcal{H}_{e}}^{\mathcal{R}_{e}}(k)\right]\\ &\cdot\left[{\bf I}_{2}\otimes{\bf R}_{\mathcal{H}_{e}}^{\mathcal{H}_{b}}(k)\right]^{\top}{\bf v}_{\mathcal{H}_{e}}^{\mathcal{H}_{b}}(k).\end{split}$ (16)
-where $\otimes$ denotes the Kronecker product, ${\boldsymbol{\Lambda}}$ is the scaling diagonal matrix, of proper size, and ${\bf R}_{\mathcal{H}_{e}}^{\mathcal{R}_{e}}(k)$ is:
-${\bf R}_{\mathcal{H}_{e}}^{\mathcal{R}_{e}}(k)={\bf R}_{\mathcal{H}_{e}}^{\mathcal{R}_{e}}{\bf R}_{\bf z}^{\mathcal{H}_{e}}(\varphi(k)).$ (17)
-We note that ${\bf R}_{\mathcal{H}_{e}}^{\mathcal{R}_{e}}$ is the fixed rotation defining the mapping between quantities in $\mathcal{H}_{e}$ and $\mathcal{R}_{e}$, and $\bf z$ is the axis of $\mathcal{H}_{e}$ mapped to the ${\bf o}$ axis of $\mathcal{R}_{e}$, according to ${\bf R}_{\mathcal{H}_{e}}^{\mathcal{R}_{e}}$.
-The ${\bf o}$ axis represents the optical axis of the camera in $\mathcal{R}_{e}$, which we assume to coincide with the ${\bf z}$ axis of $\mathcal{R}_{e}$ for negligible values of the camera tilt angle.
-Therefore, a rotation of $\varphi(k)$ about the ${\bf z}$ axis of $\mathcal{H}_{e}$ results in a viewing angle of $\varphi(k)$, as shown in Figure 5.
-[…]
-[…]
-[…]
-[…]
-Figure 5: Visualization of the rotational coupling between $\mathcal{H}_{e}$ and $\mathcal{R}_{e}$.
-The rotation matrix ${\bf R}_{\bf z}^{\mathcal{H}_{e}}(\varphi(k))$ ensures the motion correspondence between the operator’s movements and those of the FR’s TCP in the camera view, by appropriately rotating the desired velocities according to the viewing angle.
-By setting ${\bf R}_{\bf z}^{\mathcal{H}_{e}}(\varphi(k))={\bf I}_{3}$ the mapping between twists becomes viewing angle independent:
-${\bf R}_{\mathcal{H}_{e}}^{\mathcal{R}_{e}}(k)={\bf R}_{\mathcal{H}_{e}}^{\mathcal{R}_{e}},\forall k.$ (18)
+$\begin{split}{\bf v}_{\mathcal{R}_{ed}}^{\mathcal{R}_{b}}(k)=&\bigl({\bf I}_{2}\otimes{\bf R}_{\mathcal{R}_{e}}^{\mathcal{R}_{b}}(k)\bigr)\biggl\{{\boldsymbol{\Lambda}}\Bigl[\bigl({\bf I}_{2}\otimes{\bf R}_{\mathcal{H}_{e}}^{\mathcal{R}_{e}}(k)\bigr)\\ \cdot&\bigl({\bf I}_{2}\otimes{\bf R}_{\mathcal{H}_{b}}^{\mathcal{H}_{e}}(k)\bigr){\bf v}_{\mathcal{H}_{e}}^{\mathcal{H}_{b}}(k)\Bigr]\biggr\}.\end{split}$ (9)
+where $\otimes$ denotes the Kronecker product, ${\boldsymbol{\Lambda}}\in\mathbb{R}^{6\times 6}$ is the scaling diagonal matrix with ${\boldsymbol{\Lambda}}_{4:6,4:6}={\bf I}_{3}$, and
+${\bf R}_{\mathcal{H}_{e}}^{\mathcal{R}_{e}}(k)={\bf R}_{\bf z}^{\mathcal{R}_{e}}\bigl(\varphi(k)\bigr){\bf R}_{\mathcal{H}_{e}}^{\mathcal{R}_{e}}.$ (10)
+We note that ${\bf R}_{\mathcal{H}_{e}}^{\mathcal{R}_{e}}$ is the fixed rotation defining the mapping between quantities in $\mathcal{H}_{e}$ and $\mathcal{R}_{e}$, and $\bf z$ is the axis of $\mathcal{R}_{e}$, which we assume to coincide with the optical axis of the camera $\bf o$} in $\mathcal{R}_{e}$ for negligible values of the camera tilt angle.
+Therefore, a rotation of $\varphi(k)$ about the $\bf z$} axis of $\mathcal{R}_{e}$ results in a viewing angle of $\varphi(k)$.
+The $\bf z$ axis of $\mathcal{R}_{e}$ is mapped to the $\bf z$} axis of $\mathcal{H}_{e}$, according to ${\bf R}_{\mathcal{H}_{e}}^{\mathcal{R}_{e}}$.
+Hence, in the absence of compliance along rotational axes, the same behavior is achieved by rotating of $\varphi(k)$ about the $\bf z$} axis of $\mathcal{H}_{e}$, as shown in Figure 5, and (10) can be equivalently written as:
+${\bf R}_{\mathcal{H}_{e}}^{\mathcal{R}_{e}}(k)={\bf R}_{\mathcal{H}_{e}}^{\mathcal{R}_{e}}{\bf R}_{\bf z}^{\mathcal{H}_{e}}\bigl(\varphi(k)\bigr).$ (11)
+Figure 5: Visualization of the rotational coupling between $\mathcal{H}_{e}$ and $\mathcal{R}_{e}$: (top) top view; (bottom) user view; (left) $\varphi=0$; (right) $\varphi=\pi/2$.
+The rotation ${\bf R}_{\bf z}^{\mathcal{R}_{e}}\bigl(\varphi(k)\bigr)$, or equivalently ${\bf R}_{\bf z}^{\mathcal{H}_{e}}\bigl(\varphi(k)\bigr)$, ensures the motion correspondence between the operator’s movements and those of the FR’s TCP in the camera view, by appropriately rotating the desired velocities according to the viewing angle.
+By setting ${\bf R}_{\bf z}^{\mathcal{R}_{e}}(\varphi(k))={\bf I}_{3}$ (or equivalently ${\bf R}_{\bf z}^{\mathcal{H}_{e}}(\varphi(k))={\bf I}_{3}$) the mapping between twists becomes viewing angle independent:
+${\bf R}_{\mathcal{H}_{e}}^{\mathcal{R}_{e}}(k)={\bf R}_{\mathcal{H}_{e}}^{\mathcal{R}_{e}},\;\forall k\in\mathbb{N}_{0}.$ (12)
+We note that the angle $\varphi(k)$ is obtained from the exact measurements provided by the FR’s forward kinematics (or equivalently by the HD’s).
+For this reason, even during prolonged procedures, no orientation drift is accumulated, and motion correspondence is consistently preserved.
 The desired FR’s TCP homogeneous transformation ${\bf T}_{\mathcal{R}_{ed}}^{\mathcal{R}_{b}}(k)$ is obtained by numerically integrating the desired twist in (9).
-As for the desired rotation ${\bf R}_{\mathcal{R}_{ed}}^{\mathcal{R}_{b}}(k)$, it is obtained as:
+As for the desired rotation ${\bf R}_{\mathcal{R}_{ed}}^{\mathcal{R}_{b}}(k)$, it is
 ${\bf R}_{\mathcal{R}_{ed}}^{\mathcal{R}_{b}}(k)={\bf R}_{\mathcal{R}_{e}}^{\mathcal{R}_{b}}(k-1){\bf R}_{\mathcal{R}_{ed}(k)}^{\mathcal{R}_{e}(k-1)},$ (13)
-where ${\bf R}_{\mathcal{R}_{ed}(k)}^{\mathcal{R}_{e}(k-1)}$ is the rotation matrix corresponding to the axis $\frac{{\bf r}}{\left\lVert{\bf r}\right\rVert}$ and angle $\vartheta$:
-${\bf r}$ $=\left({\bf R}_{\mathcal{R}_{e}}^{\mathcal{R}_{b}}(k-1)\right)^{\top}{\boldsymbol{\omega}}_{\mathcal{R}_{ed}}^{\mathcal{R}_{b}}(k)T_{s},$ (20)
-$\vartheta$ $=\left\lVert{\bf r}\right\rVert.$ (21)
+where ${\bf R}_{\mathcal{R}_{ed}(k)}^{\mathcal{R}_{e}(k-1)}$ is the rotation matrix corresponding to the axis $\frac{{\bf r}}{\lVert{\bf r}\rVert}$ and angle $\vartheta$:
+${\bf r}={\bf R}_{\mathcal{R}_{b}}^{\mathcal{R}_{e}}(k-1)\bigl({\boldsymbol{\omega}}_{\mathcal{R}_{ed}}^{\mathcal{R}_{b}}(k)T_{s}\bigr),\quad\vartheta=\left\lVert{\bf r}\right\rVert.$ (14)
 As for the desired position ${\bf p}_{\mathcal{R}_{ed}}^{\mathcal{R}_{b}}(k)$, it is obtained as:
@@
 From ${\bf T}_{\mathcal{R}_{ed}}^{\mathcal{R}_{b}}(k)$ the desired FR’s TCP pose is obtained as:
-${\bf x}_{\mathcal{R}_{ed}}^{\mathcal{R}_{b}}(k)=\begin{bmatrix}{\bf p}_{\mathcal{R}_{ed}}^{\mathcal{R}_{b}}(k)^{\top}&{\bf q}_{\mathcal{R}_{ed}}^{\mathcal{R}_{b}}(k)^{\top}\end{bmatrix}^{\top}.$ (23)
-The eye-hand coordination mapping and scaling can be turned into an algorithmic procedure with steps provided in Algorithm 1.
+${\bf x}_{\mathcal{R}_{ed}}^{\mathcal{R}_{b}}(k)=\begin{bmatrix}{\bf p}_{\mathcal{R}_{ed}}^{\mathcal{R}_{b}}(k)^{\top},&{\bf q}_{\mathcal{R}_{ed}}^{\mathcal{R}_{b}}(k)^{\top}\end{bmatrix}^{\top}$.
+(16)
+The eye-hand coordination mapping and scaling can be turned into an algorithmic procedure as described in Algorithm 1.
 Algorithm 1 Eye-hand coordination mapping and scaling
@@
     Convert ${\bf v}_{\mathcal{H}_{e}}^{\mathcal{H}_{b}}(k)$ into ${\bf v}_{\mathcal{R}_{ed}}^{\mathcal{R}_{e}}(k)$ and rotate it according to the viewing angle using (10):
-    $\begin{split}{\bf v}_{\mathcal{R}_{ed}}^{\mathcal{R}_{e}}(k):=&\left[{\bf I}_{2}\otimes{\bf R}_{\mathcal{H}_{e}}^{\mathcal{R}_{e}}(k)\right]\left[{\bf I}_{2}\otimes{\bf R}_{\mathcal{H}_{e}}^{\mathcal{H}_{b}}(k)\right]^{\top}{\bf v}_{\mathcal{H}_{e}}^{\mathcal{H}_{b}}(k)\end{split}$
+    ${\bf v}_{\mathcal{R}_{ed}}^{\mathcal{R}_{e}}(k):=\bigl({\bf I}_{2}\otimes{\bf R}_{\mathcal{H}_{e}}^{\mathcal{R}_{e}}(k)\bigr)\bigl({\bf I}_{2}\otimes{\bf R}_{\mathcal{H}_{b}}^{\mathcal{H}_{e}}(k)\bigr){\bf v}_{\mathcal{H}_{e}}^{\mathcal{H}_{b}}(k)$
   2. 2.
     Scale ${\bf v}_{\mathcal{R}_{ed}}^{\mathcal{R}_{e}}(k)$ using $\boldsymbol{\Lambda}$ and convert it into ${\bf v}_{\mathcal{R}_{ed}}^{\mathcal{R}_{b}}(k)$:
-    ${\bf v}_{\mathcal{R}_{ed}}^{\mathcal{R}_{b}}(k)=\left[{\bf I}_{2}\otimes{\bf R}_{\mathcal{R}_{e}}^{\mathcal{R}_{b}}(k)\right]{\boldsymbol{\Lambda}}{\bf v}_{\mathcal{R}_{ed}}^{\mathcal{R}_{e}}(k)$
+    ${\bf v}_{\mathcal{R}_{ed}}^{\mathcal{R}_{b}}(k)=\bigl({\bf I}_{2}\otimes{\bf R}_{\mathcal{R}_{e}}^{\mathcal{R}_{b}}(k)\bigr)\bigl({\boldsymbol{\Lambda}}{\bf v}_{\mathcal{R}_{ed}}^{\mathcal{R}_{e}}(k)\bigr)$
 5:Compute ${\bf T}_{\mathcal{R}_{ed}}^{\mathcal{R}_{b}}(k)$ from ${\bf v}_{\mathcal{R}_{ed}}^{\mathcal{R}_{b}}(k)$ using (13) and (15)
@@
 III-B Robot interaction control
-In this section, we first formalize the contact model between the robot and the environment (Section III-B1), and subsequently, we present the free-hand strategy for limiting the interaction forces (Section III-B2).
+In this section, we first formalize the contact model between the robot and the environment (Section III-B1), and subsequently, we present the free-hand strategy for limiting the interaction forces (Section III-B2) and the other safety mechanisms embedded in the controller (Section III-B3).
 III-B1 Admittance model
-Let ${\bf x}_{\mathcal{R}_{ed}}^{\mathcal{R}_{b}}(k)=\begin{bmatrix}{\bf p}_{\mathcal{R}_{ed}}^{\mathcal{R}_{b}}(k)^{\top}&{\bf q}_{\mathcal{R}_{ed}}^{\mathcal{R}_{b}}(k)^{\top}\end{bmatrix}^{\top}$ be the desired pose and $\mathcal{R}_{c}$ the compliant frame specified by the quantities ${\bf p}_{\mathcal{R}_{c}}^{\mathcal{R}_{e}}(k)$, $\dot{\bf p}_{\mathcal{R}_{c}}^{\mathcal{R}_{e}}(k)$, $\ddot{\bf p}_{\mathcal{R}_{c}}^{\mathcal{R}_{e}}(k)$ and ${\bf q}_{\mathcal{R}_{c}}^{\mathcal{R}_{e}}(k)$.
-We do not consider compliance along the rotational axes, therefore, we assume ${\bf q}_{\mathcal{R}_{c}}^{\mathcal{R}_{e}}(k)={\bf q}_{\mathcal{R}_{ed}}^{\mathcal{R}_{e}}(k)$.
-When the FR interacts with the environment, $\mathcal{R}_{c}$ can be completely defined from ${\bf p}_{\mathcal{R}_{ed}}^{\mathcal{R}_{e}}(k)$, $\dot{\bf p}_{\mathcal{R}_{ed}}^{\mathcal{R}_{e}}(k)$ and $\ddot{\bf p}_{\mathcal{R}_{ed}}^{\mathcal{R}_{e}}(k)$, and the contact forces ${\bf f}^{\mathcal{R}_{e}}(k)\in\mathbb{R}^{3\times 1}$, by integrating the mechanical impedance equation:
-${\bf M}_{d}\ddot{\tilde{{\bf p}}}^{\mathcal{R}_{e}}(k)+{\bf K}_{D}\dot{\tilde{{\bf p}}}^{\mathcal{R}_{e}}(k)+{\bf K}_{P}\tilde{{\bf p}}^{\mathcal{R}_{e}}(k)=-\tilde{\bf f}^{\mathcal{R}_{e}}(k),$ (24)
-where ${\bf M}_{d},{\bf K}_{D},{\bf K}_{P}$ are the mass, damping, and stiffness diagonal matrices, respectively, of proper size, which are used to impose specific second-order mass-spring-damper dynamics; $\tilde{\bf p}^{\mathcal{R}_{e}}(k)={\bf p}^{\mathcal{R}_{e}}_{\mathcal{R}_{ed}}(k)-{\bf p}^{\mathcal{R}_{e}}_{\mathcal{R}_{c}}(k)$ is the position error and $\dot{\tilde{\mathbf{p}}}^{\mathcal{R}_{e}}(k)$ and $\ddot{\tilde{\mathbf{p}}}^{\mathcal{R}_{e}}(k)$ its time derivatives; $\tilde{{\bf f}}^{\mathcal{R}_{e}}(k)={\bf f}_{d}^{\mathcal{R}_{e}}(k)-{\bf f}_{e}^{\mathcal{R}_{e}}(k)$ is the force tracking error, defined as the difference between the desired force ${\bf f}_{d}^{\mathcal{R}_{e}}(k)$ and the exerted force ${\bf f}_{e}^{\mathcal{R}_{e}}(k)$, which is equal and opposite to the measured force.
-We take ${\bf f}_{d}^{\mathcal{R}_{e}}(k)={\bf 0},\forall k\in\mathbb{N}_{0}$ to make the robot behave compliantly with the environment and $\dot{\bf p}^{\mathcal{R}_{e}}_{\mathcal{R}_{ed}}(k)=\ddot{\bf p}^{\mathcal{R}_{e}}_{\mathcal{R}_{ed}}(k)={\bf 0},\forall k\in\mathbb{N}_{0}$ to reach a complete stop of the FR’s end-effector upon reaching the target position, effectively preventing any collisions caused by further movement.
-From (24), the compliant position reference to be commanded to the robot is obtained as:
-${\bf p}_{\mathcal{R}_{c}}^{\mathcal{R}_{b}}(k)={\bf p}_{\mathcal{R}_{e}}^{\mathcal{R}_{b}}(k)+{\bf R}_{\mathcal{R}_{e}}^{\mathcal{R}_{b}}(k){\bf p}_{\mathcal{R}_{c}}^{\mathcal{R}_{e}}(k).$ (25)
+Let ${\bf x}_{\mathcal{R}_{ed}}^{\mathcal{R}_{b}}(k)$ be the desired pose and $\mathcal{R}_{c}$ the compliant frame specified by the quantities ${\bf p}_{\mathcal{R}_{c}}^{\mathcal{R}_{ed}}(k)$, ${\bf q}_{\mathcal{R}_{c}}^{\mathcal{R}_{ed}}(k)$, ${\bf v}_{\mathcal{R}_{c}}^{\mathcal{R}_{ed}}(k)$, $\dot{\bf v}_{\mathcal{R}_{c}}^{\mathcal{R}_{ed}}(k)$.
+When the FR interacts with the environment, $\mathcal{R}_{c}$ can be completely defined in $\mathcal{R}_{ed}$ by integrating the mechanical impedance equation:
+${\bf M}_{d}\dot{{\bf v}}^{\mathcal{R}_{ed}}_{\mathcal{R}_{c}}(k)+{\bf K}_{D}{{\bf v}}^{\mathcal{R}_{ed}}_{\mathcal{R}_{c}}(k)+{\bf K}_{P}{\bf x}^{\mathcal{R}_{ed}}_{\mathcal{R}_{c}}(k)=-\tilde{\bf h}^{\mathcal{R}_{ed}}(k),$ (17)
+where ${\bf M}_{d},{\bf K}_{D},{\bf K}_{P}\in\mathbb{R}^{6\times 6}$ are the mass, damping, and stiffness diagonal matrices, respectively, which are used to impose specific second-order mass-spring-damper dynamics; $\tilde{{\bf h}}^{\mathcal{R}_{ed}}(k)={\bf h}_{d}^{\mathcal{R}_{ed}}(k)-{\bf h}_{e}^{\mathcal{R}_{ed}}(k)$ is the wrench tracking error, defined as the difference between the desired wrench ${\bf h}_{d}^{\mathcal{R}_{ed}}(k)$ and the exerted one ${\bf h}_{e}^{\mathcal{R}_{ed}}(k)$, which is equal and opposite to the measured wrench.
+We take ${\bf h}_{d}^{\mathcal{R}_{ed}}(k)={\bf 0},\forall k\in\mathbb{N}_{0}$ to make the robot behave compliantly with the environment.
 III-B2 Free-hand force limitation strategy
-Assuming that the environment can be approximated by a linear spring model, the exerted force at the steady state [47] is:
-${\bf f}_{ss}^{\mathcal{R}_{e}}={\bf K}_{eq}\left({\bf p}_{\mathcal{E}_{r}}^{\mathcal{R}_{e}}-{\bf p}_{\mathcal{R}_{ed}}^{\mathcal{R}_{e}}\right),$ (26)
-where ${\bf K}_{eq}=({\bf K}_{P}{\bf K}_{e})({\bf K}_{P}+{\bf K}_{e})^{-1}$ is the equivalent stiffness of the admittance stiffness ${\bf K}_{P}$ and environmental stiffness ${\bf K}_{e}$ [48], and ${\bf p}_{\mathcal{E}_{r}}^{\mathcal{R}_{e}}$ is the environment rest position.
-From (26), it follows that interaction forces can be limited through a suitable reference ${\bf p}_{\mathcal{R}_{ed}}^{\mathcal{R}_{e}}$ provided to the controller.
-Accurate limitation requires perfect knowledge of the environment that is particularly challenging to achieve, especially in a medical environment, in which the FR interacts with objects of varying stiffness and non-uniform geometry.
-Therefore, in the proposed strategy, the position reference:
-${\bf p}_{\mathcal{R}_{ed}}^{\mathcal{R}_{e}}(k)=\left[{\bf R}_{\mathcal{R}_{e}}^{\mathcal{R}_{b}}(k)\right]^{\top}{\bf p}_{\mathcal{R}_{ed}}^{\mathcal{R}_{b}}(k)-\left[{\bf R}_{\mathcal{R}_{e}}^{\mathcal{R}_{b}}(k)\right]^{\top}{\bf p}_{\mathcal{R}_{e}}^{\mathcal{R}_{b}}(k),$ (27)
+In an admittance-controlled system, the interaction wrenches emerge from the combination of the admittance parameters, the mechanical properties of the environment, and the reference provided to the controller [20].
+It follows that interaction forces can be limited through a suitable reference ${\bf x}_{\mathcal{R}_{ed}}^{\mathcal{R}_{e}}$.
+Accurate limitation requires perfect knowledge of the environment that is particularly challenging to achieve, especially in a medical environment, in which the FR interacts with surfaces of varying stiffness and non-uniform geometry.
+Therefore, in the proposed strategy, the position reference
+${\bf p}_{\mathcal{R}_{ed}}^{\mathcal{R}_{e}}(k)={\bf R}_{\mathcal{R}_{b}}^{\mathcal{R}_{e}}(k){\bf p}_{\mathcal{R}_{ed}}^{\mathcal{R}_{b}}(k)-{\bf R}_{\mathcal{R}_{b}}^{\mathcal{R}_{e}}(k){\bf p}_{\mathcal{R}_{e}}^{\mathcal{R}_{b}}(k),$ (18)
 is dynamically scaled based solely on the contact forces:
-$\underline{{\bf p}}_{\mathcal{R}_{ed}}^{\mathcal{R}_{e}}(k)=\frac{1}{1+K_{s}\left\lVert{\bf f}_{e}^{\mathcal{R}_{e}}(k)\right\rVert}{\bf p}_{\mathcal{R}_{ed}}^{\mathcal{R}_{e}}(k),$ (28)
-where $K_{s}\in\mathbb{R}^{+}$ is a proportional gain representing a trade-off between stability and responsiveness of the scaling action.
-Indeed, $K_{s}$ close to 00 reduces sudden variations of $\underline{{\bf p}}_{\mathcal{R}_{ed}}^{\mathcal{R}_{e}}(k)$, even in presence of noisy force measures, but provides less force-limiting action.
-On the other hand, $K_{s}$ significantly greater than 00 makes the scaling action more responsive, but may yield oscillations in the presence of noisy measurements.
+$\underline{{\bf p}}_{\mathcal{R}_{ed}}^{\mathcal{R}_{e}}(k)=\frac{1}{1+K_{s}\lVert{\bf f}_{e}^{\mathcal{R}_{e}}(k)\rVert}{\bf p}_{\mathcal{R}_{ed}}^{\mathcal{R}_{e}}(k),$ (19)
+where $K_{s}\in\mathbb{R}_{+}$ is a proportional gain representing a trade-off between smoothness and responsiveness of the scaling action.
+Indeed, $K_{s}$ close to 0 reduces sudden variations of $\underline{{\bf p}}_{\mathcal{R}_{ed}}^{\mathcal{R}_{e}}(k)$, even in presence of noisy force measures, but provides less force-limiting action.
+On the other hand, $K_{s}$ significantly greater than 0 makes the scaling action more responsive, but may yield oscillations in the presence of noisy measurements.
 Therefore, to increase the system’s robustness to noisy forces, exhibiting high-frequency variations, ${\bf f}_{e}^{\mathcal{R}_{e}}$ is filtered online as done for positions in (4).
@@
 However, it prevents contact forces from growing indefinitely, ensuring a limitation even in the transients of the interaction and for any orientation of the end-effector, without relying on the knowledge of the dental environment’s complex geometry.
+From a control theory perspective, it is appropriate to draw an analogy between the proposed strategy and passivity-based methods, such as TDPA [7], as both rely on modulating the reference provided to the downstream controller through an adaptive damping mechanism.
+However, a substantial difference exists in terms of control objectives.
+While passivity-based approaches provide an event-triggered modulation of the reference in response to violations of the energy balance, with the explicit goal of guaranteeing system passivity, the proposed strategy implements a continuous, non-energizing, force-driven reference modulation aimed at reducing interaction forces, without relying on explicit energy estimation.
+As a consequence, its attenuative and smoothly varying action on the reference, resulting from an appropriate choice of the gain $K_{s}$ and of the force filtering window, does not inject energy into the system, nor does it alter the underlying dynamics, in contrast to state-of-the-art variable admittance–based approaches [20, 10], which explicitly modify the controller dynamics and typically rely on explicit knowledge of the environment [48].
+As a result, the stability properties of the admittance-controlled interaction remain unaffected.
+Moreover, as discussed above, passivity-based control addresses an orthogonal objective to the proposed approach, and can therefore be combined with the eye–hand coordination and force limitation strategies to guarantee stability even in the presence of considerable communication delays.
+We consider this scenario in Section IV-B.
+III-B3 Safety mechanisms and emergency stop protocols
+From (17) and (19), the compliant pose ${\bf x}_{\mathcal{R}_{c}}^{\mathcal{R}_{b}}(k)$ to be commanded to the robot is obtained. Specifically,
+${\bf p}_{\mathcal{R}_{c}}^{\mathcal{R}_{b}}(k)$ $=\underline{{\bf p}}_{\mathcal{R}_{ed}}^{\mathcal{R}_{b}}(k)+{\bf R}_{\mathcal{R}_{ed}}^{\mathcal{R}_{b}}(k){\bf p}_{\mathcal{R}_{c}}^{\mathcal{R}_{ed}}(k),$ (20)
+${\bf q}_{\mathcal{R}_{c}}^{\mathcal{R}_{b}}(k)$ $={\bf q}_{\mathcal{R}_{ed}}^{\mathcal{R}_{b}}(k){\bf q}_{\mathcal{R}_{c}}^{\mathcal{R}_{ed}}(k),$ (21)
+where ${\bf x}^{\mathcal{R}_{ed}}_{\mathcal{R}_{c}}$, taking the semantics of a pose error, is obtained by time-integrating $\dot{\bf v}^{\mathcal{R}_{ed}}_{\mathcal{R}_{c}}$ from (17) twice.
+In addition to its compliant interaction properties and the embedded force limitation strategy, the controller incorporates two safety mechanisms and implements an emergency stop protocol.
+Regarding the safety mechanisms, the first ensures the stability of the FR in the absence of updated references from the EHCC ${\bf x}_{\mathcal{R}_{ed}}^{\mathcal{R}_{b}}(k)$, such as in the event of a communication loss, by continuing to track the last valid one ${\bf x}_{\mathcal{R}_{ed}}^{\mathcal{R}_{b}}(k-1)$.
+The second mechanism prevents abrupt motions of the FR by commanding ${\bf x}_{\mathcal{R}_{e}}^{\mathcal{R}_{b}}(k)$ whenever the generated command ${\bf x}_{\mathcal{R}_{c}}^{\mathcal{R}_{b}}(k)$ deviates excessively, either in translation or orientation, based on translational and rotational safety thresholds that can be configured by the operator according to the task requirements.
+With regard to the emergency stop, it focuses on the monitoring of the interaction forces.
+Whenever $\lVert{\bf f}^{\mathcal{R}_{e}}\rVert(k)$ exceeds a predefined safety threshold $f_{th}$, the controller forwards ${\bf x}_{\mathcal{R}_{e}}^{\mathcal{R}_{b}}(k)$ as the new reference while continuing to compute the admittance control law.
+In this way, abrupt stopping is avoided, while still allowing the interaction forces to decrease progressively.
 III-C Haptic feedback controller
@@
 In impedance systems, the interaction is captured by both the contact force ${\bf f}_{e}^{\mathcal{R}_{e}}$ and the position error $\tilde{\bf p}^{\mathcal{R}_{e}}$, the latter being characterized by slower transients than the former, which exhibits a step dynamics.
-As proposed in [49], replacing ${\bf f}_{e}^{\mathcal{R}_{e}}$ with $\tilde{\bf p}^{\mathcal{R}_{e}}$ to generate the force feedback makes each contact state transparent to the operator, preventing them from having to deal with sudden force changes they are not used to from previous training and fieldworks.
+As proposed in [41], replacing ${\bf f}_{e}^{\mathcal{R}_{e}}$ with $\tilde{\bf p}^{\mathcal{R}_{e}}={\bf p}^{\mathcal{R}_{e}}_{\mathcal{R}_{ed}}-{\bf p}_{\mathcal{R}_{c}}^{\mathcal{R}_{e}}$ to generate the force feedback makes each contact state transparent to the operator, preventing them from having to deal with sudden force changes they are not used to from previous training and fieldworks.
 Therefore, the proposed HFC defines a second (virtual) impedance model, generating the force feedback from the motion error of the admittance controller:
 ${\bf f}_{h}^{\mathcal{H}_{b}}(k)={\bf K}_{h}\tilde{{\bf p}}^{\mathcal{H}_{b}}(k)+{\bf D}_{h}\dot{\tilde{{\bf p}}}^{\mathcal{H}_{b}}(k),$ (22)
-where ${\bf K}_{h}$ and ${\bf D}_{h}$ respectively are virtual stiffness and damping diagonal matrices of the proper size.
-The position error $\tilde{{\bf p}}^{\mathcal{H}_{b}}(k)$ and its time derivative $\dot{\tilde{{\bf p}}}^{\mathcal{H}_{b}}(k)$ are defined, respectively, as:
-$\tilde{{\bf p}}^{\mathcal{H}_{b}}(k)$ $={\bf R}_{\mathcal{H}_{e}}^{\mathcal{H}_{b}}(k)\left[{\bf R}_{\mathcal{H}_{e}}^{\mathcal{R}_{e}}(k)\right]^{\top}\tilde{{\bf p}}^{\mathcal{R}_{e}}(k),$ (30)
-$\dot{\tilde{{\bf p}}}^{\mathcal{H}_{b}}(k)$ $={\bf R}_{\mathcal{H}_{e}}^{\mathcal{H}_{b}}(k)\left[{\bf R}_{\mathcal{H}_{e}}^{\mathcal{R}_{e}}(k)\right]^{\top}\dot{\tilde{{\bf p}}}^{\mathcal{R}_{e}}(k).$ (31)
-We note that, as in (17), the rotation matrix ${\bf R}_{\bf z}^{\mathcal{H}_{e}}(\varphi(k))$ is used to provide the force feedback correspondence with the interaction observed in the camera view.
-With regard to the ${\bf K}_{h}$ and ${\bf D}_{h}$ parameters of the virtual system in (29), they are tuned to provide a natural force feedback to the human operator while the FR interacts with the environment, decoupling it from the admittance control dynamics.
-Specifically, the stiffness matrix ${\bf K}_{h}$ can be tuned to achieve the desired transient response, while the damping matrix ${\bf D}_{h}$ can be tuned to mitigate vibrations.
-We envisage the possibility that the generated force feedback may exceed the range of forces that the HD device can exert.
-Therefore, the HFC saturates the norm of the force feedback $\lVert{\bf f}_{h}^{\mathcal{H}_{b}}(k)\rVert$ to the maximum exertable force $f_{MAX}^{\mathcal{H}_{b}}\in\mathbb{R}^{+}$:
-$\underline{\mathbf{f}}_{h}^{\mathcal{H}_{b}}(k)=\begin{cases}\mathbf{f}_{h}^{\mathcal{H}_{b}}(k),&\text{if}~{}\lVert\mathbf{f}_{h}^{\mathcal{H}_{b}}(k)\rVert<f_{MAX}^{\mathcal{H}_{b}};\\ f_{MAX}^{\mathcal{H}_{b}}\dfrac{\mathbf{f}_{h}^{\mathcal{H}_{b}}(k)}{\lVert\mathbf{f}_{h}^{\mathcal{H}_{b}}(k)\rVert},&\text{otherwise}.\end{cases}$ (32)
-Finally, given the slow dynamics of $\tilde{{\bf p}}^{\mathcal{R}_{e}}$, the human operator may perceive residual force feedback even when the robot is no longer interacting.
-To address this, we introduce a force dead-band $f_{db}\in\mathbb{R}^{+}$ to prevent the HD from reacting to such residuals:
-$\underline{{\bf f}}_{h}^{\mathcal{H}_{b}}(k)=\begin{cases}\underline{{\bf f}}_{h}^{\mathcal{H}_{b}}(k),&\text{if}~{}\left\lVert{\bf f}_{e}^{\mathcal{R}_{e}}(k)\right\rVert>f_{db};\\ {\bf 0},&\text{otherwise}.\end{cases}$ (33)
-IV Experimental validation
-To evaluate the performance of the proposed FH-HBTS, we conduct interaction experiments using a phantom.
-Each experiment involves reaching a designated target tooth with an inactive dental scaler and applying a force to its surface, keeping the tool stationary.
-During the experiments, the robot remains out of the human operator’s view, who relies solely on the visual feedback from the camera and the force feedback from the HD.
-We select three target teeth: the lower right first premolar, the upper right central incisor, and the upper right second premolar, as shown in Figure 6.
-The chosen targets are significant since each requires a different tool orientation, making the use of eye-hand coordination essential.
-For each target, we perform three interaction experiments with active force limitation and three without.
-With the aim of confirming the impact of the proposed strategy on the safety of interaction, we estimate, from the data recorded during the experiments, the conditional expectation of the interaction forces given a virtual penetration commanded by the operator.
-The conducted analysis is reported in Section IV-B.
-Figure 6: Target teeth for dental scaling experiments: 1) lower right first premolar; 2) upper right central incisor; 3) upper right second premolar.
-In order to quantitatively assess the operation of the eye-hand coordination, we conduct a free-space experiment which concerns commanding the same spatial motion in the HD space, with different viewing angles.
-The results of the task execution are reported in Section IV-C.
-Finally, we provide an experimental demonstration of the end-to-end system by conducting interaction experiments with the three target teeth in both scenarios, this time emulating the dental scaling procedure.
-The outcomes of a dental scaling procedure execution, representative of the overall results are shown in Section IV-D.
+where ${\bf K}_{h}\in\mathbb{R}^{3\times 3}$ and ${\bf D}_{h}\in\mathbb{R}^{3\times 3}$ respectively are virtual stiffness and damping diagonal matrices, where
+$\begin{bmatrix}\tilde{{\bf p}}^{\mathcal{H}_{b}}(k)\\ \dot{\tilde{{\bf p}}}^{\mathcal{H}_{b}}(k)\end{bmatrix}=\Bigl({\bf I}_{2}\otimes\bigl({\bf R}_{\mathcal{H}_{e}}^{\mathcal{H}_{b}}(k){\bf R}_{\mathcal{R}_{e}}^{\mathcal{H}_{e}}(k)\bigr)\Bigr)\begin{bmatrix}\tilde{{\bf p}}^{\mathcal{R}_{e}}(k)\\ \dot{\tilde{{\bf p}}}^{\mathcal{R}_{e}}(k)\end{bmatrix}$.
+(23)
+We note that, as in (9), the rotation matrix ${\bf R}_{\bf z}^{\mathcal{R}_{e}}(k)$ (or equivalently ${\bf R}_{\bf z}^{\mathcal{H}_{e}}(k)$) in ${\bf R}_{\mathcal{H}_{e}}^{\mathcal{R}_{e}}(k)$ is used to provide the force feedback correspondence with the interaction observed in the camera view.
+This approach is particularly suited to the considered application, as it guarantees stability while providing the operator with a natural and intuitive tactile experience that closely resembles manual execution, especially in light of Remark 2.
+Within this framework, the parameters ${\bf K}_{h}$ and ${\bf D}_{h}$ of the virtual system in (22) are tuned to shape the perceived interaction, with the stiffness matrix ${\bf K}_{h}$ governing the transient response and the damping matrix ${\bf D}_{h}$ contributing to vibration attenuation.
+Moreover, the use of a second virtual impedance model to generate the force feedback moderates the influence of transient force spikes associated with sudden contacts, while preserving informative and natural haptic feedback.
+We envisage the possibility that the generated feedback may exceed the range of forces that the HD device can exert.
+Therefore, the HFC saturates the norm of the force feedback $\lVert{\bf f}_{h}^{\mathcal{H}_{b}}(k)\rVert$ to the maximum exertable force $f_{\text{max}}^{\mathcal{H}_{b}}\in\mathbb{R}_{+}$:
+$\underline{\mathbf{f}}_{h}^{\mathcal{H}_{b}}(k)=\begin{cases}\mathbf{f}_{h}^{\mathcal{H}_{b}}(k),&\text{if}~\lVert\mathbf{f}_{h}^{\mathcal{H}_{b}}(k)\rVert<f_{\text{max}}^{\mathcal{H}_{b}};\\ f_{\text{max}}^{\mathcal{H}_{b}}\dfrac{\mathbf{f}_{h}^{\mathcal{H}_{b}}(k)}{\lVert\mathbf{f}_{h}^{\mathcal{H}_{b}}(k)\rVert},&\text{otherwise}.\end{cases}$ (24)
+Finally, given the slow dynamics of $\tilde{{\bf p}}^{\mathcal{R}_{e}}$, the operator may perceive residual force even when the robot is no longer interacting.
+To address this, we introduce a force dead-band $f_{db}\in\mathbb{R}_{+}$ to prevent the HD from reacting to such residuals:
+$\overline{{\bf f}}_{h}^{\mathcal{H}_{b}}(k)=\begin{cases}\underline{{\bf f}}_{h}^{\mathcal{H}_{b}}(k),&\text{if}~\bigl\lVert{\bf f}_{e}^{\mathcal{R}_{e}}(k)\bigr\rVert>f_{db};\\ {\bf 0},&\text{otherwise}.\end{cases}$ (25)
+III-D Integrated mechanical end-effector
+In order to successfully perform any free-hand dental procedure, the operator relies on three basic factors: (a) the procedure-specific clinical tool; (b) the surgical field view, which may require external illumination and tools to show obstructed or poorly lit areas, e.g., a mirror; (c) the sense of touch to ensure the right pressure to be exerted on the area of interest, in order to achieve a safe and effective procedure.
+In this work, we design an integrated mechanical end-effector that can provide all the tools and sensory capabilities required to perform free-hand dental procedures.
+Figure 6: Integrated mechanical end-effector for free-hand procedures. Visible in the left side of the figure is also one of the dental models used in the experiments.
+With reference to Figure 6, the designed end-effector consists of five main components: (1) a 3D-printed mechanical interface connecting the robot’s flange with the force/torque sensor; (2) a force/torque sensor measuring contact forces; (3) a 3D-printed tool socket supporting the clinical tool and the endoscopic camera: the interface consists of a rigid structure to enable proper transmission of forces from the tool to the sensor, a circular hole, adjustable via a bolt-and-nut mechanism, to accommodate tools of different sizes, and a movable camera-support ring, allowing for adjustments of the view and light beam angle; (4) an high definition endoscopic camera, equipped with LED illumination to ensure good visibility even in poorly lit areas; (5) a clinical tool for performing the procedure of interest: the tool is placed in the supporting socket after being wrapped with friction tape to ensure good mechanical compliance, emulating the skin friction between the operator’s hand and the tool; the tool used for the chosen use case is a scaler.
+Although the current implementation does not fulfill yet all clinical requirements, such a design is compatible with future clinical adoption, as the end-effector is inherently compatible with standard sterilization and infection-control procedures commonly adopted in dentistry.
+Indeed, clinical instruments are intrinsically designed to withstand autoclave sterilization, and the literature reports the availability of sterilizable endoscopic cameras [9], as well as autoclavable or low-temperature–sterilizable materials, including medical-grade 3D-printed polymers, which would enable full sterilization of the entire assembly [50].
+Moreover, the tool socket has been designed with smooth external surfaces and a reduced number of crevices, thereby facilitating cleaning and reprocessing.
+In addition, the socket enables rapid tool replacement by simple extraction and insertion of the clinical instrument, without requiring disassembly of the end-effector or disconnection of sensing and vision components.
+Components upstream of the end-effector, such as the force/torque sensor and the robot flange, which are not compatible with autoclave sterilization, can be isolated using sterile drapes or disposable sterile barriers, as commonly adopted in medical robotic systems [60].
+With respect to aerosol management, a critical concern in dental procedures such as SRP, the proposed design does not interfere with conventional mitigation strategies.
+In a clinical deployment, aerosol generation could be managed through standard high-volume evacuation systems and saliva ejectors, without requiring any modification to the design.
+IV Technical validation
+To evaluate the performance of the proposed FH-HBTS, we conduct a set of experiments including static force application (Section IV-B), dynamic dental scaling motions (provided in the supplementary materials), and a free-space experiment (Section IV-C).
+During the experiments, the robot is out of the operator’s view, who relies solely on the visual feedback from the camera and the force feedback from the HD.
+We select three target teeth: the lower right first premolar, the upper right central incisor, and the upper right second premolar.
+The chosen targets are significant as each requires a different tool orientation, making the use of eye-hand coordination essential.
 IV-A Experimental setup
-The experimental setup involves the following equipment and materials:
-  • a leader control station featuring a 6-degrees-of-freedom (DOF) force-feedback HD [50] and a workstation.
-The latter is a general purpose computer executing the leader controller loop, and communicating with the FR over an Ethernet network.
-In such a network, delays and packet losses are negligible, as is the case of an healthcare facility;
-  • a FR consisting of the 6-DOF UR10 robotic arm by Universal Robots [51].
-It is equipped with an integrated mechanical end-effector for free-hand dental procedures.
-Its design is discussed in Section IV-A1;
-  • a dental phantom, i.e. 3D-printed life-sized model of a human mouth.
-IV-A1 Integrated mechanical end-effector
-In order to successfully perform any free-hand dental procedure, the operator relies on three basic factors: (a) the procedure-specific clinical tool; (b) the surgical field view, which may require external illumination and tools to show obstructed or poorly lit areas e.g., a mirror; (c) the sense of touch to ensure the right pressure to be exerted on the area of interest, in order to achieve a safe and effective procedure.
-In this work, we design an integrated mechanical end-effector, shown in Figure 7, that can provide all the tools and sensory capabilities required to perform free-hand dental procedures.
-The designed end-effector consists of five main components:
-  1. a)
-    a 3D-printed mechanical interface connecting the robot’s flange with the force/torque sensor;
-  2. b)
-    a force/torque sensor measuring contact forces. The sensor used is the FT150 6-axis force/torque sensor by Robotiq [52];
-  3. c)
-    a 3D-printed tool socket supporting the clinical tool and the endoscopic camera.
-The interface consists of a rigid structure to enable proper transmission of forces from the tool to the sensor, a circular hole, adjustable via a bolt-and-nut mechanism, to accommodate tools of different sizes, and a movable camera-support ring, allowing for adjustments of the view and light beam angle;
-  4. d)
-    an HD endoscopic camera, equipped with LED illumination to ensure good visibility even in poorly lit areas;
-  5. e)
-    a clinical tool for performing the procedure of interest.
-The tool is placed in the supporting socket after being wrapped with friction tape to ensure good mechanical compliance, emulating the skin friction between the operator’s hand and the tool.
-The tool used for the chosen use case is a scaler.
-Figure 7: Integrated mechanical end-effector for free-hand procedures, consisting of: 1) 3D-printed mechanical interface connecting the robot’s flange with the force/torque sensor; 2) force/torque sensor; 3) 3D-printed tool socket supporting the clinical tool and the endoscopic camera; 4) HD endoscopic camera; 5) clinical tool.
-Visible in the left side of the figure is also the dental phantom used.
-IV-A2 Eye-hand coordination controller parametrization
-[…]
-With regard to the scaling factors, we set $\beta=1$ and ${\boldsymbol{\Lambda}}=diag(0.8,0.8,0.8,1.0,1.0,1.0)$ to achieve high control accuracy on the translational axes without significantly reducing the workspace size.
-IV-A3 Admittance controller parametrization
-Since the interaction experiments require modest compliance and high damping along the normal axis, and greater and more precise positioning along the tangential axes, we set ${\bf K}_{P}=\text{diag}(50000,50000,6000)$, ${\bf K}_{D}=\text{diag}(2500,2500,12000)$ and ${\bf M}_{d}=\text{diag}(50,50,50)$.
-Regarding the force limitation strategy, we identify a suitable trade-off between responsiveness and stability by setting $K_{s}=0.06$, and by filtering the exerted forces with a moving average filter of window size $n=500$.
-IV-A4 Virtual feedback parametrization
-With regard to the virtualization of the force feedback, the virtual impedance matrices are tuned to provide the human operator with a natural sensation, ensuring both smooth transient response and reduced vibrations.
-Therefore, we set ${\bf K}_{h}=\text{diag}(0.1,0.1,0.1)$, ${\bf D}_{h}=\text{diag}(0.5,0.5,0.5)$.
-[…]
+The experimental setup involves the following equipment and materials: (I) a leader control station featuring a 6-degrees-of-freedom (DOF) force-feedback HD [^2] https://www.3dsystems.com/haptics-devices/touch and a workstation: the latter is a general purpose computer executing the leader controller loop, and communicating with the FR over an Ethernet network, where delays and packet losses are negligible; (II) a FR consisting of the 6-DOF UR10 arm [^3] https://www.universal-robots.com/, equipped with the integrated mechanical end-effector of Section III-D, where the force/torque sensor is the Robotiq FT150444https://robotiq.com; (III) a high-strength resin dental model and a 3D-printed dental phantom, both life-sized models of a human mouth.
+The former is used for the experiments described in this section, while the latter only appears in the accompanying video.
 IV-B Force limitation assessment
-[…]
-Given the conditional probability function $p\left(a(k)|b(k)\right)$, we estimate the conditional expectation $E\left(A|B\right)$, and the conditional variance $\text{Var}(A|B)$ following [55].
-In Figure 8 are reported the conditional expectations of the exerted force $E\left(f_{e}^{\mathcal{R}_{e},\perp}\middle|\left\lVert{\bf p}_{\mathcal{R}_{ed}}^{\mathcal{R}_{e}}\right\rVert\right)$ and the conditional standard deviation, obtained as the square root of the conditional variance $\text{Var}\left(f_{e}^{\mathcal{R}_{e},\perp}\middle|\left\lVert{\bf p}_{\mathcal{R}_{ed}}^{\mathcal{R}_{e}}\right\rVert\right)$, in both scenarios.
-Visual inspection showcases a significant reduction in the expected normal interaction force in Scenario A, given the same virtual penetration commanded.
-Figure 8: Comparison between the conditional expectation of the interaction force in Scenario A and Scenario B.
-To evaluate the effectiveness of the force limitation strategy, we compare $E\left(f_{e}^{\mathcal{R}_{e},\perp}\middle|\left\lVert{\bf p}_{\mathcal{R}_{ed}}^{\mathcal{R}_{e}}\right\rVert\right)$ and $\text{Var}\left(f_{e}^{\mathcal{R}_{e},\perp}\middle|\left\lVert{\bf p}_{\mathcal{R}_{ed}}^{\mathcal{R}_{e}}\right\rVert\right)$ in both scenarios.
-To test whether the conditional variances differ significantly, we perform Levene’s test and Bartlett’s test.
-Levene’s test shows that the conditional variances are significantly different (p-value $p\approx 0$), indicating that force limitation strategy affects the dispersion of normal force $f_{e}^{\mathcal{R}_{e},\perp}(k)$ with respect to virtual penetration $\left\lVert{\bf p}_{\mathcal{R}_{ed}}^{\mathcal{R}_{e}}(k)\right\rVert$.
-Bartlett’s test, which is more sensitive to normal distributions, confirms this observation ($p\approx 0$).
-This suggests that the difference in variance is not only due to the sampling effect but is a structural feature of the collected data.
-We compare the conditional expectation using a t-test for independent samples.
+To assess the effectiveness of the proposed force limitation strategy, we perform $9$ interaction experiments, with an average duration of $\sim$20\text{\,}\mathrm{s}$\$ each, 3 for each target tooth, without force limitation (Scenario A), $9$ more with force limitation (Scenario B) and $9$ with a passivity-based architecture (Scenario C), specifically TDPA-HD, where admittance control replaces Cartesian impedance control adopted in [42].
+In the latter scenario, the force limitation is intentionally kept disabled in order to isolate the effects of TDPA-HD.
+For each scenario, we record the data sequence $\bigl\{a(k),b(k)\bigr\}_{k=1}^{N}$, sampling at a frequency of $125\text{\,}\mathrm{H}\mathrm{z}$, where we set the force norm $\left\lVert{\bf f}_{e}^{\mathcal{R}_{e}}(k)\right\rVert$ as $a(k)\in\mathcal{A}\subseteq\mathbb{R}_{+}$, and the commanded virtual penetration $\bigl\lVert{\bf p}_{\mathcal{R}_{ed}}^{\mathcal{R}_{e}}(k)\bigr\rVert$ as $b(k)\in\mathcal{B}=\{b\in\mathbb{R}|b_{\text{min}}=$0\text{\,}\mathrm{m}$\leq b\leq b_{\text{max}}=$0.007\text{\,}\mathrm{m}$\$\}.
+We discretize $\mathcal{B}$ with the step $b_{\text{step}}=$0.0001\text{\,}\mathrm{m}$\$, selected based on the repeatability of the FR, while $b_{\text{max}}$ is the maximum virtual penetration observed across all experiments.
+By binning data in $\mathcal{B}$, following [56], we estimate the conditional expectation of the exerted force $E\bigl(\lVert{\bf f}_{e}^{\mathcal{R}_{e}}\rVert\big|\lVert{\bf p}_{\mathcal{R}_{ed}}^{\mathcal{R}_{e}}\rVert\bigr)$ and the conditional standard deviation, obtained as the square root of the conditional variance $\text{Var}\bigl(\lVert{\bf f}_{e}^{\mathcal{R}_{e}}\rVert\big|\lVert{\bf p}_{\mathcal{R}_{ed}}^{\mathcal{R}_{e}}\rVert\bigr)$, in all three scenarios.
+Visual inspection of Figure 7 showcases a significant reduction in the expected interaction force in Scenario B, given the same commanded virtual penetration.
+Notably, TDPA-HD [42] (Scenario C) cannot reduce the interaction force in quasi-static contacts due to the energy being approximately zero.
+Figure 7: Comparison between the conditional expectation of the interaction force in all three scenarios.
+To evaluate the effectiveness of the force limitation strategy, we compare $E\bigl(\lVert{\bf f}_{e}^{\mathcal{R}_{e}}\rVert\big|\lVert{\bf p}_{\mathcal{R}_{ed}}^{\mathcal{R}_{e}}\rVert\bigr)$ and $\text{Var}\bigl(\lVert{\bf f}_{e}^{\mathcal{R}_{e}}\rVert\big|\lVert{\bf p}_{\mathcal{R}_{ed}}^{\mathcal{R}_{e}}\rVert\bigr)$ in all three scenarios.
+We assess differences in conditional variances through Levene’s and Bartlett’s tests in the comparisons involving Scenario B, i.e., B–A and B–C.
+Both tests indicate significantly different variances ($p\approx 0$), showing that the force limitation strategy affects the dispersion of the force norm $\bigl\lVert{\bf f}_{e}^{\mathcal{R}_{e}}(k)\bigr\rVert$ with respect to virtual penetration $\bigl\lVert{\bf p}_{\mathcal{R}{ed}}^{\mathcal{R}_{e}}(k)\bigr\rVert$.
 Since Levene’s test suggests a significant variance difference, we apply Welch’s t-test, which is more robust in the case of unequal variances.
-The t-test returned a t-statistic value of $t=-6.18$, confirming that the normal force is lower in Scenario A, and $p\approx 0$.
-Since such value is below the significance threshold $\alpha=0.05$, we can reject the null hypothesis and conclude that the conditional expectations of normal force in the two scenarios are significantly different.
+In the B–A comparison, the test yields a t-statistic t=4.52t=-4.52 with $p\approx 0$, confirming a statistically significant force reduction in Scenario B.
+Beyond statistical significance, the magnitude of this effect is notable: the comparison yields a Cohen’s d=0.63d=-0.63, indicating a significant force reduction when the force limitation strategy is enabled.
+The corresponding confidence interval [0.91,0.36][-0.91,-0.36] does not include zero, further supporting the robustness and consistency of the observed effect across the explored range of virtual penetration.
+An analogous analysis is conducted for the B-C comparison.
+Also in this case, Welch’s t-test confirms a statistically significant reduction of the normal force with respect to Scenario C ($p\approx 0$), indicating that the force limitation strategy yields lower contact forces than the passivity-based control architecture.
+The observed effect is comparable in magnitude to that of the B-A comparison (Cohen’s d=0.68d=-0.68 and confidence interval [0.97,0.41][-0.97,-0.41]), further supporting the robustness of the proposed approach across different control architectures.
 IV-C Eye-hand coordination assessment
-In order to perform a formal assessment of eye-hand coordination, we execute a teleoperated task in the free space.
-The human operator performs a translatory motion with the HD, simultaneously rotating the viewing angle by rotating the HD’s stylus by an angle $\varphi$ about the $\bf z$ axis of $\mathcal{H}_{e}$.
+To perform a formal assessment of eye-hand coordination, we execute a teleoperated task in the free space.
+The operator performs a translatory motion with the HD, simultaneously rotating the viewing angle by rotating the HD’s stylus by an angle $\varphi$ about the $\bf z$ axis of $\mathcal{H}_{e}$.
 The validation is considered completed if the motion transferred to the robot varies as the viewing angle does.
 Figure 8 shows the $\varphi$ angle, the position of HD’s stylus and the position of the FR, respectively.
-(a) $\varphi$ angle about the $z$ axis of $\mathcal{H}_{e}$.
-(b) Position of HD’s stylus.
-(c) Position of the FR.
 Figure 8: Eye-hand coordination assessment.
-[…]
+Results showcase that, as long as the angle $\varphi=$0\text{\,}\mathrm{rad}$\$ (from $0\text{\,}\mathrm{s}$ to $10\text{\,}\mathrm{s}$), a motion along the $x$ axis of $\mathcal{H}_{b}$ is mapped into a motion along the $x$ axis of $\mathcal{R}_{b}$.
+When the angle turns into $\varphi=-\pi/$2\text{\,}\mathrm{rad}$\$, a motion along the $x$ axis of $\mathcal{H}_{b}$ is mapped into a motion along the $z$ axis of $\mathcal{R}_{b}$ (from $20\text{\,}\mathrm{s}$ to $30\text{\,}\mathrm{s}$).
 These results confirm that the mapping of motion between HD space and FR space is affected by the rotation of the visual field.
 However, they do not provide any suggestion regarding the effectiveness of eye-hand coordination for the human operator.
-Therefore, in the accompanying video [^1] This paper has supplementary downloadable material available at http://ieeexplore.ieee.org, provided by the authors.
-This includes a multimedia MOV format movie clip, showing the end-to-end system experimental demonstration.
-This material is 51.4 MB in size.
-we provide a comprehensive qualitative assessment of the latter.
-IV-D End-to-end system experimental demonstration
-We provide a demonstration of the end-to-end system’s operation in the accompanying video, where we showcase the execution of the dental scaling procedure on the three target teeth in Scenario A.
-In this section, we report the outcomes of the execution of a dental scaling procedure in Scenario A, highlighting two instants of the experiment (Figure 10), and compare them against those obtained in Scenario B (Figure 11).
-To avoid repetition, the analysis focuses on the outcomes of executing a dental scaling action on a single target tooth, the upper right second premolar.
-In Figure 10d and Figure 11b, the normal forces exerted by the FR in both scenarios are shown.
-A visual inspection showcases a significant reduction in interaction forces when the free-hand force limitation strategy is active, with a decrease of approximately $50\%$ in Scenario A, thereby confirming the validation presented in Section IV-B.
-The effectiveness of this strategy is made even more evident by comparing Figure 10c and Figure 11a.
-Notably, in Scenario A, where the interaction forces are lower, the virtual penetration commanded to the controller is even greater than in Scenario B.
-Nevertheless, as shown in Figure 10c, the implementation of this strategy enables the scaling of the provided references, effectively reducing the contact forces and enhancing the system’s safety.
-[…]
-[…]
-(c) Norm of the admittance controller reference
-(d) Contact forces
-(e) Forces rendered on the HD
-Figure 10: Snapshots and outcomes of the dental scaling experiment in Scenario A.
-(a) Norm of the admittance controller reference
-(b) Contact forces
-(c) Forces rendered on the HD
-Figure 11: Outcomes of the dental scaling experiment in Scenario B.
-Finally, in Figure 10e and Figure 11c the forces rendered on the HD in both scenarios are shown.
-As expected, the results are consistent between the two scenarios, while both showing high similarity to exerted force profiles as well as slow dynamics.
-Such aspects enhance naturalness for the human operator.
-V Discussions and conclusions
+Therefore, in the accompanying video we provide a comprehensive qualitative assessment of the latter.
+V In-vitro pre-clinical validation
+To validate the feasibility of the FH-HBTS in performing free-hand dental procedures, we conduct an in-vitro experimental trial replicating a SRP procedure under more representative clinical conditions.
+The procedure is carried out on the dental model, coated with an acrylic polymer–based suspension containing titanium dioxide to simulate artificial calculus, using a fully functional ultrasonic scaler activated via a foot pedal at the leader control station.
+A total of 20 participants with varying levels of experience in the dental field are recruited: (Group 1) dental students who had performed at least 20 in-vivo dental scaling procedures $(31.6\%)$; (Group 2) dentists with less than 5 years of professional experience $(36.8\%)$; (Group 3) dentists with more than 5 years of professional experience $(31.6\%)$.
+Before the experimental phase, all participants complete a structured training session aimed at ensuring competent use of the system, during which they practice approaching the dental surface, maintaining stable contact, and performing preliminary calculus removal.
+After training, participants perform artificial calculus removal on selected tooth surfaces: lower left central incisor, lower left second molar, lower right first molar, relying on both visual and haptic feedback.
+For each trial, each surface is uniformly coated with a thin layer of artificial calculus.
+The procedure is considered successful if the participant is satisfied with the outcome and none of the safety mechanisms described in Section III-B3 are triggered.
+A single representative trial, illustrating the overall experimental procedure, is provided in the accompanying video.
+V-A Data collection and analysis
+To comprehensively evaluate system performance, operator experience, and operational safety, a set of objective and subjective outcome measures is collected and analyzed.
+These include indicators of clinical efficacy, task efficiency, interaction safety, and perceived workload and usability.
+Differences across participant experience levels are assessed for all outcome measures using Kruskal–Wallis tests.
+System efficacy is evaluated by an independent expert dentist, who assesses calculus removal using a six-point ordinal scale ranging from no removal to complete removal.
+Task efficiency is quantified through the median execution time per dental surface, while interaction safety is assessed by analyzing the magnitude of the interaction forces exerted during the procedure.
+Both metrics are compared with reference values reported in the literature for manual execution.
+Finally, system usability is evaluated through subjective measures, including the NASA-Task Load Index (NASA-TLX) questionnaire for perceived workload, the System Usability Scale (SUS) questionnaire for usability, and additional 5-point Likert-scale questions to further characterize the operator interaction.
+The complete questionnaire is reported in the supplementary materials.
+V-B Results
+All participants were able to successfully complete the entire experimental procedure.
+Regarding system efficacy, the mean cleanness score per participant, averaged over the three teeth, ranges from a minimum of $3.66$ to a maximum of $4.28$, indicating a generally high and homogeneous level of task execution.
+No statistically significant differences between groups are observed, as confirmed by the Kruskal–Wallis test conducted at a significance level of $p=0.05$.
+Regarding execution time, the duration of the procedure per tooth showcases substantial variability, ranging from $85\text{\,}\mathrm{s}$ to $212\text{\,}\mathrm{s}$.
+The median execution times are $153.5\text{\,}\mathrm{s}$ for Group 1, $162.0\text{\,}\mathrm{s}$ for Group 2, and $147.5\text{\,}\mathrm{s}$ for Group 3.
+Reference values reported in [49] for manual performance of the same procedure indicate a median execution time of $115.5\text{\,}\mathrm{s}$.
+Although the teleoperated procedure is therefore slower than manual execution, participants exhibited an average reduction of $20.4\%$ in completion time in the second trial, followed by a further $23.2\%$ reduction in the third trial.
+This trend reflects a clear learning-curve effect, indicating that operators progressively familiarize themselves with the system and improve their ability to control the remote tool.
+This suggests that, with more extensive training, teleoperated performance could gradually approach manual execution levels.
+In terms of interaction forces, during the experimental trial, the force magnitude results in an overall mean of $1.78\pm 0.19\,\mathrm{N}$ for Group 1, $1.86\pm 0.34\,\mathrm{N}$ for Group 2, and $1.96\pm 0.32\,\mathrm{N}$ for Group 3, with average maximum peaks of $9.44\pm 3.09\,\mathrm{N}$, $9.80\pm 4.7\,\mathrm{N}$, and $11.47\pm 4.41\,\mathrm{N}$, respectively.
+When comparing groups, Kruskal-Wallis tests find no statistically significant differences between experience levels.
+The comparison with reference values reported in the literature for dentists performing the procedure free-hand highlights a clear discrepancy.
+In particular, [57] reports an average interaction force of $0.28\pm$0.33\text{\,}\mathrm{N}$\$, with average maximum peaks of approximately $0.60\text{\,}\mathrm{N}$.
+These values are close to the tissue safety thresholds reported in the literature for dental procedures, commonly identified at around $0.5\text{\,}\mathrm{N}$ [17].
+Notably, these reference values are also compatible with those measured in a dedicated free-hand SCP procedure we conduct on the dental model mounted on the force/torque sensor to directly measure interaction forces.
+Further details on this experiment are provided in the supplementary materials.
+This difference reflects an inherent limitation of the system, which is primarily attributable to the mechanical characteristics of the adopted industrial robot and to the resolution of the force/torque sensor, not specifically designed to accurately manage the low-magnitude forces typical of dental interactions.
+Despite the higher measured forces, no superficial damage was observed on the dental model, suggesting that the system remains safe under the adopted experimental conditions.
+However, these results should be regarded as scalable, as the interaction forces are strongly influenced by the employed hardware setup and are expected to decrease when adopting an engineered system tailored to the target application, particularly in terms of mechanical compliance and force/torque sensing resolution.
+The results of the Kruskal–Wallis tests on questionnaire responses are reported in Table I.
+No statistically significant differences emerge across any NASA-TLX dimension, SUS score, or 5-point Likert item ($p>0.05$), indicating that subjective evaluations are consistent across experience levels.
+Across all groups, the NASA-TLX results indicate a consistently moderate perceived workload, with contained mental, physical, and temporal demands, limited variability in performance ratings, and generally low effort and frustration.
+Overall, the task is perceived as manageable and not overly demanding.
+System usability is evaluated positively, with SUS scores falling within the good-usability range [8], confirming that the system is considered intuitive and easy to use.
+Likert-scale evaluations further highlight high ratings for interaction-related aspects, including motion and force correspondence, first contact and prolonged interaction perception, and the naturalness of the haptic feedback (mean $\in[3.43,4.33]$), suggesting coherent and supportive sensory feedback.
+Ergonomics is also positively evaluated (mean $\in[3.71,4.17]$), exceeding the manual reference condition of $3$ and indicating improved ergonomic and postural comfort during teleoperation.
+Fidelity achieves only moderately positive scores, likely due to limitations in depth perception (mean $\in[1.71,2.33]$), preventing accurate situational awareness.
+Group 1 Group 2 Group 3 K–W test Mean $\pm$ STD Mean $\pm$ STD Mean $\pm$ STD $\chi^{2}$ p NASA-TLX Mental Demand $30.00\pm 27.57$ $51.43\pm 15.74$ $36.67\pm 23.38$ 2.91 0.23 Physical Demand $26.67\pm 24.22$ $20.00\pm 11.55$ $16.67\pm 19.66$ 0.70 0.70 Temporal Demand $36.67\pm 23.38$ $42.86\pm 13.80$ $36.67\pm 15.06$ 0.52 0.77 Performance $43.33\pm 8.17$ $7.14\pm 13.80 33$$3.33\pm 10.33$ 2.42 0.30 Effort $16.67\pm 19.66$ $8.57\pm 10.69$ $23.33\pm 15.06$ 2.79 0.25 Frustration $30.00\pm 27.57$ $28.57\pm 27.95$ $33.33\pm 16.33$ 0.11 0.94 Overall $26.19\pm 18.65$ $26.94\pm 13.36$ $25.71\pm 14.26$ 2.00 0.37 SUS $77.92\pm 13.46$ $81.07\pm 9.11$ $78.75\pm 10.34$ 0.51 0.78 5-point Likert Scale Ergonomics $3.83\pm 0.75$ $3.71\pm 0.76$ $4.17\pm 0.98$ 0.95 0.62 Fidelity $3.17\pm 1.33$ $00\pm 0.58 3.3$$33\pm 0.52$ 0.87 0.65 Motion correspondence $4.17\pm 0.41$ $4.00\pm 0.82$ $4.33\pm 0.82$ 0.78 0.68 Force correspondence $4.00\pm 0.89$ $4.14\pm 0.38$ $4.17\pm 0.75$ 0.17 0.92 Depth perception $2.17\pm 0.98$ $1.71\pm 0.76$ $2.33\pm 0.52$ 2.19 0.33 Interaction events $3.83\pm 0.98$ $3.43\pm 0.98$ $4.00\pm 1.10$ 1.30 0.52 Interaction effects $3.83\pm 0.98$ $3.43\pm 0.79$ $4.17\pm 0.75$ 2.73 0.26 Feedback naturalness $4.00\pm 0.89$ $3.71\pm 0.95$ $3.83\pm 1.17$ 0.33 0.85
+TABLE I: Summary of the NASTA-TLX and 5-point Likert scale questionnaire responses and SUS scores.
+VI Conclusions and Future Perspectives
 This paper proposed a FH-HBTS for free-hand dental procedures.
@@
 We also improved the accuracy of the procedure by introducing human tremor filtering and dynamic scaling of velocities.
-By providing the eye-hand coordination and the TCP control, we enabled the human operator to perceive the motion correspondence between their movements and those of the end-effector, observed in the camera view, preserving manual dexterity and expertise.
+Through eye-hand coordination and TCP control, we enabled the operator to perceive the motion correspondence between their movements and those of the end-effector, observed in the camera view, preserving manual dexterity and expertise.
 Additionally, by appropriately transforming the virtualized force feedback, we also ensured the force correspondence, enabling the human operator to experience a highly natural sensory perception.
-Finally, we enabled safe interactions, by scaling the motion references provided to the admittance controller.
+Finally, we enabled safe interactions by scaling the motion references provided to the admittance controller.
 The proposed strategy allows for accurate force limitation in any contact state and for any orientation of the interaction tool, without requiring any prior modeling of the intervention site.
 We believe that our FH-HBTS represents a crucial step towards a broader adoption of robotics in dentistry, as it allows for accurate treatments, while preserving the operator’s cognitive and technical abilities.
-With respect to manual procedures, our system allows to preserve operator’s ergonomics, track operator’s performance, and enable remote healthcare.
-We emphasize that the proposed system allows to collect haptic and motion data.
-While this feature is crucial during the training phase of new operators for the evaluation of learning metrics, i.e learning curves [56], it is also central towards the deployment of data-driven control architectures that leverage demonstrations to synthesize control policies [57].
+With respect to manual procedures, our system allows improving operator’s ergonomics, tracking operator’s performance, and perspectively enabling remote healthcare.
+We emphasize that the proposed system allows collecting haptic and motion data.
+While this feature is crucial during the training phase of new operators for the evaluation of learning metrics, i.e learning curves [71], it is also central towards the deployment of data-driven control architectures that leverage demonstrations to synthesize autonomous control policies [19].
 Finally, these data can also be used to learn cost functions [21], and hence intentions, of experts fulfilling a given task.
 In the context of dental procedures, this means that data collected from experts executing the procedure can be used to train policies for autonomous robots.
-With our future work we plan to extend the FH-HBTS with autonomous patient motion compensation to further enhance the system’s safety, and to benchmark the performance against the da Vinci Surgical SystemTM.
-Also, we plan to conduct experiments with professionals and active dental tools both to confirm the system usability under more representative clinical conditions and to collect expert policies for (semi-)autonomous procedures.
+With our future work we plan to extend the FH-HBTS with autonomous patient motion compensation to further enhance system’s safety and felxibility, and to benchmark the performance against the da Vinci Surgical SystemTM.
+In addition, future experimental testbeds will be designed to progressively incorporate clinically relevant sources of complexity that are not yet addressed in the present study, including constrained inter-proximal access, interaction with soft tissues, fluid presence, and limited workspace conditions.
+In this context, we plan to collect expert demonstrations for the development of (semi-)autonomous dental procedures.
+Beyond algorithmic extensions, a key direction of future work concerns the evolution of the experimental platform itself, which will be redesigned towards a fully engineered, application-driven system, explicitly accounting for the mechanical, sensing, and ergonomic requirements of dental procedures, and addressing the hardware-related limitations identified in the present study.
 Acknowledgments
-The authors would like to thank Sabatino Panico for contributing to the system calibration and experimental campaigns.
+The authors would like to thank X for their assistance during the in-vitro pre-clinical validation and Y for supporting the EHCC implementation and end-to-end system validation, making this work possible.